### v1

# PEEB: Part-based Image Classifiers with an Explainable and Editable Language Bottleneck

Abstract
CLIP-based classifiers rely on the prompt containing a {class name} that is known to the text encoder. That is, CLIP performs poorly on new classes or the classes whose names rarely appear on the Internet (e.g., scientific names of birds). For fine-grained classification, we propose PEEB – an explainable and editable classifier to (1) express the class name into a set of pre-defined text descriptors that describe the visual parts of that class; and (2) match the embeddings of the detected parts to their textual descriptors in each class to compute a logit score for classification. In a zero-shot setting where the class names are unknown, PEEB outperforms CLIP by a large margin (10 in accuracy). Compared to part-based classifiers, PEEB is not only the state-of-the-art on the supervised-learning setting (88.80% accuracy) but also the first to enable users to edit the class definitions to form a new classifier without re-training. Compared to concept bottleneck models, PEEB is also the state-of-the-art in both zero-shot and supervised learning settings.
PEEB: Part-based Image Classifiers with an
Explainable and Editable Language Bottleneck
Thang M. Pham††thanks: Equal contribution thangpham@auburn.edu Peijie Chen* pzc0018@auburn.edu Tin Nguyen* ttn0011@auburn.edu
Seunghyun Yoon syoon@adobe.com Trung Bui bui@adobe.com Anh Nguyen anh.ng8@gmail.com
Auburn University Adobe Research
1 Introduction
Fine-grained bird classification (Wah et al., 2011; Van Horn et al., 2015) is a long-standing challenge in computer vision. Yet, state-of-the-art (SOTA) bird classifiers often have one or more of the following four limitations. First, many models both CNN-based (Krause et al., 2016) and ViT-based (He et al., 2022) are inherently black-box. That is, they have no built-in mechanisms that explain to users how a decision is made, e.g., which bird traits make a model think a given bird is Indigo Bunting? Second, many bird classifiers claim to be explainable (Chen et al., 2019; Donnelly et al., 2022) by comparing the input image with a set of learned, part-based prototypes or natural language concepts (Fig. 1a–b). Yet, such prototypes are feature vectors and therefore not editable by users. Third, textual concept-based models operate at the image level (Fig. 1a) and it is unknown what image details match a given descriptor Menon and Vondrick (2023); Yang et al. (2023). Fourth, most classifiers require either training-set images in a supervised-learning setting or demonstration images in a zero-shot setting (Xian et al., 2018; Zhu et al., 2018). This requirement is impractical when building a classifier for a novel class whose photos do not yet exist in the database.
To address the above problems, we propose PEEB, an image classifier that is both explainable and editable via natural language. PEEB classifies images based only on the textual descriptor of bird parts provided by humans or GPT-4 (no images needed) and grounds the descriptors to the visual bird parts for more fine-grained explanations (Fig. 1c). While PEEB leverages CLIP’s encoders (Radford et al., 2021), it uses no class names (e.g., Indigo Bunting) in the prompt. In contrast, CLIP-based classifiers (Radford et al., 2021) and its extensions (Pratt et al., 2022; Menon and Vondrick, 2023) rely so heavily on the known class names in the prompt that their accuracy drops significantly when the names are removed or replaced by uncommon alters (Sec. 5.1).
We first define parts of interest that humans typically observe to recognize a bird. We follow CUB annotations to have 15 parts and later reduce to 12 by merging the directional parts (e.g. left wing and right wing are combined to wings) (Appendix C). Using GPT-4, we construct a textual descriptor (OpenAI, 2023) to describe each bird part of every species (Appendix C). Next, PEEB generates 12 visual part embeddings, then localizes the 12 bird parts in the input image based on these embeddings (Fig. 2). The unnormalized distance (logits) between the input image and every class would be the sum of the 12 dot products between the paired visual and textual part embeddings (Fig. 3). Besides being editable by humans, PEEB outperforms concept-based explainable classifiers across different zero-shot settings and prototypical part networks in a supervised fashion.
To our knowledge, all existing public bird-image datasets (listed in Table A4) are limited in size (less than 100K images per dataset) and in the number of classes (less than 1,500 species per dataset), impeding large-scale, vision-language, contrastive learning research. Therefore, for our pre-training, we build Bird-11K, an unprecedentedly large bird-image dataset of 290K images and 11K species, i.e., basically all bird species on Earth (Sec. 3). Bird-11K is constructed from 7 existing bird datasets and 55K new images that we collect from the Macaulay Library.
Our main findings are:111Code and dataset construction scripts are released on https://github.com/anguyen8/peeb.
- 1.
2.
Our part-level pre-trained PEEB outperforms CLIP-based classifiers by +8 to +29 points in bird classification across CUB, NABirds, and iNaturalist datasets (Sec. 5.2).
3.
PEEB allows defining new classes in text during the test time without re-training the models (Fig. 2). Besides interpretability and editability, PEEB outperforms text concept-based methods in the generalized-zero-shot setting (Sec. 5.3) and the zero-shot setting by Elhoseiny et al. (2017) (Sec. 5.4), showing strong generalization capabilities.
4.
When compared with explainable classifiers in supervised learning on CUB, PEEB scores an 88.80% accuracy, which is competitive to the best CUB classifiers trained using supervised learning (81–87% accuracy) in the literature (Sec. 5.5) that are often not editable.
5.
PEEB is potentially applicable to other domains such as dogs, cats, fish or butterflies. In the dog domain, PEEB outperforms explainable methods while achieving comparable performance to SOTA black-box classifiers (Sec. 5.6).
2 Related Work
Standard CNNs and Transformers It is common to build bird classifiers based on standard CNNs such as ResNets (He et al., 2016) or ViTs (He et al., 2022). Although high-performing, these models do not admit an inherent explanation interface (Gunning et al., 2021) and therefore rely on post-hoc interpretability methods, which tend to offer inaccurate and unstable, after-the-fact explanations (Rudin, 2019; Bansal et al., 2020). In our work, the textual part descriptors form a natural-language bottleneck interface that enables users to observe and edit the bird attributes that contribute to each final prediction. That is, users can re-program the classifier without having to re-train any network (see Fig. 2).
Prototypical Part Networks There are bird classifiers, such as ProtoPNet (Chen et al., 2019), designed with an explainability objective to learn prototypes i.e., learnable concepts (Nauta et al., 2021; Donnelly et al., 2022; Nguyen et al., 2022; Nauta et al., 2022; Kim et al., 2022a; Wang et al., 2021). Yet, because such prototypes are real-valued vectors, it is unknown how much users could interpret and use them in a downstream task. In contrast, PEEB is a bird classifier that relies on part-based concepts but allows users to define the textual descriptors of the birds of interest, while prior prototype-based classifiers require complete re-training if any prototype needs modifications.
Textual-based Concept Bottlenecks Recent vision-language models (VLMs) are often considered interpretable due to their reliance on natural language concepts. However, several works (Samuel et al., 2021; Yuksekgonul et al., 2023; Esfandiarpoor and Bach, 2023), which rely on specific conditions such as differences in class-wise captions Esfandiarpoor and Bach (2023) or learned concept weights Yang et al. (2023); Panousis et al. (2023); Oikarinen et al. (2023), are constrained and lack the capability for generalization. Another line of work with diverse approaches to process textual concepts for training typically lacks immediate editability and often requires re-training or additional steps for new concepts (Ji et al., 2018; Zhu et al., 2018; Paz-Argaman et al., 2020; Chen et al., 2020; Kousha and Brubaker, 2021; Rao et al., 2023; Han et al., 2023). CLIP-based classifiers rely heavily on having correct class names rather than descriptors in text prompts (Pratt et al., 2022; Menon and Vondrick, 2023) and thus are neither explainable nor editable to humans. Unlike CLIP-based models, PEEB reveals what image details are being used for classification by matching descriptors and visual bird part (e.g. beak in Fig. 3).
Attribute-based Classifiers The Attribute Label Embedding (ALE) approach (Akata et al., 2015) employs a fixed set of attributes and trains an attribute-to-label weight matrix for zero-shot classification. Several studies (Xu et al., 2020; Hanouti and Le Borgne, 2023) highlight its effectiveness on datasets like CUB, SUN (Xiao et al., 2010), and AWA (Xian et al., 2019). However, ALE has a limitation in terms of flexibility i.e. expanding label numbers necessitates overhauling the attribute set, weight matrix, and model re-training. This rigidity contrasts with our target for a scalable and adaptable framework, leading us to explore alternatives beyond ALE despite its success.
3 Bird-11K Dataset
3.1 Dataset construction
We combine bird images from 7 distinct datasets with 55K images (10,534 classes) collected from Cornell’s Macaulay Library, to form a unified Bird-11K dataset 222We do not redistribute the published datasets but release a script to reconstruct Bird-11K. (Sec. D.1) for large-scale pre-training. To the best of our knowledge, Bird-11K, comprising 440,934 images spanning 11,183 classes, is the first bird dataset that encompasses almost all species on Earth. Since PEEB learns to match visual parts to textual descriptors, it requires that bird images be distinctly visible and sufficiently large for accurate part localization and matching (See Sec. E.3 for ablation study). However, small and “hard-to-see” bird images in Bird-11K make the dataset noisy and the training complex. Thus, we employ OWL-ViT (Minderer et al., 2022) to detect bird objects in all images using the query “bird” and filter out images with the detected bird’s bounding boxes smaller than pixels. The mean Intersection over Union (IoU) between the predicted bird boxes and ground-truth boxes on CUB dataset is 0.91.
As class names from different datasets are either general (e.g. Cardinal) or specific (e.g. Yellow Cardinal, Northern Cardinal), we retain only the specific species for more diverse training and exclude all general classes to circumvent class ambiguity. Following these filtration steps, the refined Bird-11K dataset retains 294,528 images across 10,811 classes (Table A4). For each species in Bird-11K, we generate a set of part-based descriptors using GPT-4. Details of the descriptor generation are provided in Appendix C.
3.2 Dataset splits for contrastive pre-training
Two distinct zero-shot settings emerged in the recent literature. The first setting is conventional zero-shot (ZSL), which ensures a model is not exposed to any test classes during training. In our experiments with Bird-11K, we ensure test classes from datasets like CUB or NABirds remain unseen during pre-training. For example, when testing on CUB, we exclude all CUB classes and their images, and similarly for NABirds, to prevent any prior exposure to the model. For evaluation, we employ the traditional CUB split proposed by Akata et al. (2015) and two harder splits: Super-Category-Shared/Exclusive (SCS/SCE) by Elhoseiny et al. (2017). For example, in ZSL on CUB, we exclude all CUB classes in Bird-11K for pre-training and fine-tune only on the corresponding training set given by a ZSL split.
The second setting is generalized zero-shot (GZSL), where models like CLIP are trained on large-scale datasets that may inadvertently include a subset of the testing classes or images. In the GZSL test, we exclude all test sets from CUB, NABirds, and iNaturalist, and directly evaluate the pre-trained model without further fine-tuning.
4 Method
4.1 Backbone: OWL-ViT bird-part detector
OWL-ViT (Minderer et al., 2022) is an open-vocabulary object detector that detects objects in an image, given text queries, even if those objects are unseen during training. OWL-ViT consists of four components: a standard Vision Transformer image encoder, and an architecturally identical text encoder, a box regression head, and a box classification head. While the box regression head is a three-layer Multilayer Perceptron (MLP) followed by a GELU activation (Hendrycks and Gimpel, 2016) for the first two linear layers, the box classification head is simply a linear layer to project the visual embeddings to the same dimensional space with text embeddings.
4.2 Part-based, explainable, and editable bird classifier (PEEB)
Architecture The PEEB model is composed of two encoders—an image encoder and a text encoder—as well as three key components: a linear projection block, a part MLP (Multi-Layer Perceptron), and a box MLP. Part MLP layer is designed to map visual part embeddings to the same dimensional space with descriptor embeddings for computing logits that can be directly used for classification ( in Fig. 3). This design allows PEEB to perform arbitrary ways of classification. The remaining components of the model are adopted from the OWL-ViT framework. Details of the components are provided in Appendix A.
Model Inference For a given image, we first employ the 12 part names as textual queries and select the visual part embeddings based on the cosine similarity. These selected visual part embeddings are then simultaneously fed into both the part MLP and the box MLP. The box MLP will predict the bounding box coordinates of each part. The output from the part MLP is matched against the encoded text descriptors, with predictions being made based on their aggregated cosine similarity. Specifically, we compute the dot product between the selected visual part embeddings and text embeddings of part descriptors, summing up the resulting twelve scores to form a logit for each class. Predictions are then obtained by applying the to these logits. An overview diagram of PEEB’s inference process can be found in Fig. 3.
4.3 Training strategy
We empirically find that solely training the part MLP layer does not achieve the desired classification accuracy, prompting us to update the image encoder. However, re-training this encoder impacts the linear projection and box MLP layers. As a result, we have to train all components together. All the components are initialized from the corresponding components in OWL-ViT, except for our proposed part MLP. Our training strategy has two phases: two-stage pre-training on the large-scale Bird-11K dataset and fine-tuning on downstream tasks. We provide full hyperparameter details of PEEB in Sec. A.8.
Objectives There are three objectives to train PEEB: (a) Train the part MLP layer contrastively to maximize the similarity between related part-descriptor pairs while minimizing the unrelated pairs using symmetric cross-entropy (CE) loss (Radford et al., 2021); (b) Train the linear projection layer to mimic OWL-ViT’s behaviors (i.e. the similarity matrix) for part selection with symmetric CE loss; and (c) Train the box MLP layer for bounding box regression with DETR losses (Zheng et al., 2020) i.e. a linear combination of corner-to-corner distance loss and GIoU loss (Rezatofighi et al., 2019). We provide detailed loss functions in Sec. A.10.
Challenges One problem emerges when we jointly train all components together: the model learns at a significantly slow pace since PEEB needs to learn to optimize two symmetric cross-entropy losses while maintaining the high-quality predicted boxes. To tackle this problem, we split the pre-training phase into two stages: (1) train the image encoder and part MLP layer with the first objective; then (2) train the linear projection and box MLP layers with the second and third objectives to accordingly adjust their weights to the changes in the image encoder. Notably, the text encoder is always frozen as it was designed for open vocabulary, its generalizability to unseen texts (i.e., descriptors of an unseen bird) should be preserved.
4.3.1 Pre-training on Bird-11K dataset
Stage 1, contrastive learning with teacher model: The image encoder and part MLP layer are jointly trained using the symmetric CE loss, which is particularly suitable for PEEB as it learns the mapping between visual parts and descriptors. In this stage, we follow the teacher model – OWL-ViT to select part embeddings to ensure the selected part embeddings are semantically meaningful, e.g., can be used for box prediction (Fig. A2).
Stage 2, eliminate teacher dependency:
As the image encoder is modified in stage 1, the linear project and the box MLP needs to be updated accordingly. Specifically, we consider the teacher model OWL-ViT as ground-truths and train the linear projection layer using symmetric CE loss (Fig. A3, 1a–c, 2a–c). For box MLP, given the absence of human-annotated boxes for individual parts, we obtain pseudo labels sourced from OWL-ViT as ground-truths for the training with DETR losses (Fig. A3, 2d). In this training step, the image encoder is frozen while the part MLP layer is not involved. After two-step training, PEEB can perform zero-shot classification while providing the mappings between visual parts and descriptors as faithful explanations.
4.3.2 Fine-tuning on target datasets
We can further fine-tune the pre-trained model on downstream tasks, e.g., CUB, NABirds and iNaturalist to compare with other baselines (e.g. prototype-based approaches). In this phase, to adapt to the downstream tasks, all components except the text encoder are trained jointly and the loss function for part MLP is changed from symmetric CE to CE while other losses are kept intact.
5 Experiments & Results
We conduct systematic experiments to compare the generalization ability of PEEB with explainable methods on two zero-shot settings: GZSL (Sec. 5.1 – 5.3) and ZSL (Sec. 5.4). Notably, our method demonstrates significantly superior performance in GZSL and generalizes better in ZSL setting.
Moreover, following Donnelly et al., 2022, we finetune and evaluate PEEB on downstream tasks to measure the transferability compared to other part- and text concept-based methods (Sec. 5.5). Results show that PEEB achieves decent performance while providing more faithful explanations (Figs. 4, A10, A11 and A12). According to the evaluation for visual part localization (Sec. E.7) and qualitative analysis (Appendix G), we find that the box prediction performance of PEEB is comparable with OWL-ViT.
Original Descriptors (a)Incorrect Descriptors (b)
Blue JayBlue Jay 0.0059Blue Jay 0.0058
M&V
Blue Jay 0.6899 (c)Least Tern 0.0611 (d)
PEEB
5.1 CLIP-based classifiers depend mostly on class names (not part descriptors)
M&V shows that incorporating descriptors generated by GPT-3 (Brown et al., 2020) to class names increases CLIP accuracy on downstream tasks. Yet, it remains unknown how useful the descriptors are compared to the randomized descriptors for M&V’s method.
Experiment To address the concern, we conduct two systematic studies: (1) randomly swap a set of part descriptors among classes in CUB, NABirds, and iNaturalist datasets and measure the contribution of descriptors by the difference in model performance on CUB test set (i.e., 200-way classification) using original and random descriptions. (2) Replace the common names with scientific names in CUB, NABirds, and iNaturalist and measure the difference in model performance. The first experiment aims to study the contributions of descriptions to the CLIP-based method, and the second experiment tries to understand the dependencies of the class names.
Results When random descriptors are used, M&V ’s accuracy drops marginally by -0.9 points, while PEEB’s performance significantly deteriorates, indicating a strong dependence on accurate descriptions. However, when only descriptors are used without class names (Table 1) or class names are replaced by scientific names (Table 2), M&V ’s accuracy reduces drastically to nearly random chance, underscoring the reliance of CLIP-based classifiers on class names.
5.2 Part-based pre-trained PEEB outperforms CLIP-based classifiers
The dependence on class names suggests CLIP is potentially exposed to class names during training. Thus, we compare PEEB with the CLIP-based classifiers in the GZSL setting for a fair comparison.
Experiments We conduct two-stage pre-training for PEEB on Bird-11K where CUB, NABirds, and iNaturalist test sets are excluded for evaluation. Note that the contrastive pre-training is on part level, that is, PEEB does not have direct access to the class labels.
Results PEEB outperforms the baselines across all three datasets, achieving improvements of (+10 to +12 points), (+28 to +29 points) and (+8 to +9 points) on CUB, NABirds and iNaturalist, respectively (Table 2).
| CLIP | M&V | PEEB (ours) | ||
| Using class names | ✗ | ✗ | ||
| Original Descriptors | 52.02 | 53.78 | 5.89 | 64.33 |
| Incorrect Descriptors | n/a | 52.88 | 0.59 | 0.88 |
5.3 PEEB is superior to text concept-based classifiers on GZSL setting
The advancements in Large Language Models (LLMs) led to the emergence of text concept-based approaches (Yuksekgonul et al., 2023; Yan et al., 2023; Esfandiarpoor and Bach, 2023; Panousis et al., 2023) aiming to provide a more flexible and explainable classifier. We compare PEEB in the GZSL setup on CUB test set with text concept-based methods.
Experiment We conduct an experiment with the same setting as in Sec. 5.2. While PEEB does not use specific concepts, the descriptors it utilizes (described in Appendix C) function as open-vocabulary concepts. Most text concept-based methods in GZSL require pre-defined concept sets, limiting their applicability to similar datasets like NABirds or iNaturalist. Hence, our evaluation is focused solely on the CUB dataset.
Results PEEB exhibits superior GZSL performance, outperforming recent text concept-based approaches by +3 to +10 points (Table 3). Unlike the traditional approaches, which primarily rely on predefined concept pools, PEEB utilizes domain-specific, natural language descriptors. This distinctive approach enables PEEB to adapt more flexibly to various datasets such as NABirds and iNaturalist. In contrast, concept pool-based methods often necessitate manual concept updates and potential re-training for such extensions.
| Method | CUB | Concept-type |
| FuDD (2023) | 54.30 | Class-wise Captions |
| Han et al. (2023) | 56.13 | Evolving Descriptions |
| PCBM (2023) | 61.00 | Pre-defined Concept Pool |
| Yan et al. (2023) | 60.27* | Pre-defined Concept Pool |
| LaBo (2023) | 54.19 | Pre-defined Concept Pool |
| PEEB (Ours) | 64.33 | Class-wise Descriptors |
5.4 PEEB generalizes to traditional ZSL
Some vision-language approaches enable the model to process certain textual-based concepts Ji et al. (2018); Zhu et al. (2018); Paz-Argaman et al. (2020); Chen et al. (2020); Kousha and Brubaker (2021); Rao et al. (2023), and therefore, ZSL is used to measure their generalization capability. In this evaluation, we employed the traditional ZSL split on CUB from Akata et al. and the Super-Category-Similar/Exclusive (SCS/SCE) proposed by Elhoseiny et al. on CUB and NABirds. The SCS (easy) and SCE (hard) split are designed based on the species hierarchy to intentionally have two levels of difficulties from the ZSL test.
Experiment Adhere to the conventional ZSL setting; we exclude all the CUB or NABirds classes for pre-training and fine-tune the model following Akata et al. or Elhoseiny et al.’s split. We randomly select 10% from the training set as the validation set and choose the checkpoints based on the lowest validation loss.
Results PEEB outperforms all baselines across 3 test splits (from CUB and NABirds) by (+4 to +10 points) in terms of harmonic mean, indicating that PEEB is more generalized to not only seen classes (80.78 vs 65.80) but also unseen classes (all other results in Table 4). The easy tests (SCS) guarantee the presence of classes similar to (but distinct from) the ones in the training set. Therefore, all baselines that learn better on the training set tend to have better accuracy in the easy test. Conversely, the hard splits (SCE) ensure the test classes are from different categories of the training classes. This distinction makes the hard test a more accurate metric for assessing a model’s generalization ability. PEEB excels over all baselines by (+5 to +15 points) (accuracy) for SCE split and +2.64 points (accuracy) compared to CLORE.
Furthermore, we also evaluate the model that is pre-trained without all CUB and NABirds classes and compare them with CLIP and M&V methods on CUB and NABirds. Interestingly, PEEB outperforms both baselines with (+10 to +12) points on CUB and (+1 to +3) points on NABirds (Sec. E.1). This finding further substantiates the generalization capability of our method.
| Methods | CUB | NABirds | ||||
| Seen | Unseen | Harmonic | Seen | Unseen | Harmonic | |
| Data split by Akata et al. (2015) | ||||||
| CLORE (2022) | 65.80 | 39.10 | 49.05 | n/a | ||
| PEEB (ours) | 80.78 | 41.74 | 55.04 | |||
| SCS/SCE splits by Elhoseiny et al. (2017) | ||||||
| SCS | SCE | Harmonic | SCS | SCE | Harmonic | |
| (Easy) | (Hard) | (Easy) | (Hard) | |||
| SGA-DET (2018) | 42.90 | 10.90 | 17.38 | 39.40 | 9.70 | 15.56 |
| GRZSL (2018) | 44.08 | 14.46 | 21.77 | 36.36 | 9.04 | 14.48 |
| ZEST (2020) | 48.57 | 15.26 | 23.22 | 38.51 | 10.23 | 16.17 |
| CANZSL (2020) | 45.80 | 14.30 | 21.12 | 38.10 | 8.90 | 14.43 |
| DGRZSL (2021) | 45.48 | 14.29 | 21.75 | 37.62 | 8.91 | 14.41 |
| DPZSL (2023) | 45.40 | 15.50 | 23.11 | 40.80 | 8.20 | 13.66 |
| PEEB (ours) | 44.66 | 20.31 | 27.92 | 28.26 | 24.34 | 26.15 |
5.5 Finetuning pre-trained PEEB on CUB-200 yields a competitive explainable classifier in supervised learning
We compare our model with explainable methods, such as ProtoPNet Chen et al. (2019), using the conventional pre-training and fine-tuning approach on Bird-11K and the CUB dataset. We focus on evaluating how effectively the pre-trained PEEB transfers to downstream tasks, providing insights into its adaptability and performance.
Experiment We further fine-tune all components of the pre-trained PEEB on the CUB dataset for comparison with other explainable methods. We fine-tune the pre-trained model for 30 epochs and select the best checkpoints based on validation loss. We provide the detailed hyperparameters in Table A2.
Results In the CUB dataset, PEEB trained from OWL-ViT backbone achieves 77.80% accuracy. With pre-training on Bird-11K, our model sets a new standard: 86.73% and 88.80% accuracy with two different backbones (Table 5), exceeding Deformable ProtoPNet (86.4%) and even ProtoTree (87.20%). PEEB distinctly maps descriptors to visual inputs, facilitating easier debugging and clearer understanding as demonstrated in Fig. 4.
| Methods | Model size | Backbone | Accuracy |
| Base (ViT) (2021) | 22M | DeiT-S (2021) | 84.28 |
| – Concept bottleneck classifiers | |||
| Concept Bottleneck Models (2020) | 11M | ResNet-18 (2016) | 80.10 |
| CPM (2023) | 155M | ViT-B/16 (2021) | 72.00 |
| CDM (2023) | 155M | ViT-B/16 | 74.31 |
| LaBo (2023) | 427M | ViT-L/14 | 81.90 |
| – Part-based classifiers | |||
| ProtoPNet (2019) | 22M | DeiT-S | 84.04 |
| ProtoTree (2021) | 92M* | ResNet-50 (2016) | 87.20* |
| TesNet (2021) | 79M | Dense121 (2017) | 84.80 |
| Deformable ProtoPNet (2022) | 23M | ResNet-50 | 86.40 |
| ProtoPFormer (2022) | 22M | DeiT-S | 84.85 |
| PEEB (ours) – w/o pre-training | 155M | OWL-ViT | 77.80 |
| PEEB (ours) | 155M | OWL-ViT | 86.73 |
| PEEB (ours) | 155M | OWL-ViT | 88.80 |
5.6 Applying PEEB to Dog classification
PEEB can be applied across various domains, including dogs, cats, fish, and butterflies. We specifically demonstrate its applicability in the dog domain in this section.
Data preparation
We first define parts of interest that humans typically observe to distinguish dog breeds. Following PartImageNet He et al. (2021), we have 4 parts of interest: head, body, legs and tail and add two more parts: muzzle and ears as they also play a key role in dog recognition. Then, we prompt GPT-4 to get the descriptors for 6 parts per dog breed for 140 classes from Dog-140 – a large-scale dataset we construct from ImageNet-21K for pre-training (Sec. D.2). Finally, we run OWL-ViT model to detect the corresponding boxes for 6 pre-defined parts.
Experiment
Following bird experiments, we pre-train PEEB on Dog-140 and then further fine-tune on Stanford Dogs for evaluation. The experimental setup allows us to evaluate the pre-trained PEEB in GZSL setting and the fine-tuned PEEB on the downstream task. Additionally, we fine-tune PEEB from OWL-ViT backbone to get the lower bound performance on Stanford Dogs.
Results
PEEB achieves 74.17% accuracy on Stanford Dogs without pre-training on a large-scale dataset, which is similar to PEEB performance on CUB with 77.80% (Table 5). Gathering more images from ImageNet-21K for pre-training enhances PEEB’s accuracy to 87.38% for the pre-trained PEEB and 92.20% for fine-tuned PEEB (Table 6), while providing fine-grained explanations (Fig. 5). Interestingly, PEEB not only outperforms XAI methods but also achieves performance comparable to SOTA black boxes (92.20% vs 97.00%).
| Methods | Model size | Backbone | Accuracy |
| – Explainable methods | |||
| FCAN (2016) | 50M | ResNet-50 | 84.20 |
| RA-CNN (2017) | 144M | VGG-19 | 87.30 |
| ProtoPNet (2019) | 22M | DeiT-S | 77.30 |
| Deformable ProtoPNet (2022) | 23M | ResNet-50 | 86.50 |
| – SOTA black-box classifiers | |||
| TransFG (2022) | 86M | ViT-B/16 | 92.30 |
| ViT-Net (2022b) | 86M | DeiT-B | 93.60 |
| SR-GNN (2022) | 32M | Xception | 97.00 |
| PEEB (ours) | |||
| + trained on Stanford Dogs | 155M | OWL-ViT | 74.17 |
| + pre-trained on Dog-140 | 155M | OWL-ViT | 87.37 |
| + pre-trained on Dog-140 & | 155M | OWL-ViT | 92.20 |
| fine-tuned on Stanford Dogs | |||
6 Discussion and Conclusion
Explainability and editability PEEB stands out as a transparent and editable classifier to users by grounding the text descriptors to the visual parts in the image (Fig. 4-bottom). This transparent decision-making process plays an important role in user understanding. For instance, in Fig. 4 (upper right), the challenge of understanding why the model predicts accurately persists, particularly when we already know that the descriptors are incorrect. In contrast, PEEB’s explanations not only make errors like the mismatch between throat and wings more apparent but also enable users to adjust descriptions, thereby improving model accuracy without the need for retraining (Fig. 2).
We introduce PEEB, a pioneering part-based, explainable, and editable bird classifier that leverages textual descriptors for bird parts. By grounding natural language descriptors with visual features, PEEB brings transparency to its decision-making. Besides, PEEB achieves superior performance in both GZSL and ZSL settings compared to existing state-of-the-art explainable models. It underscores PEEB’s robustness and versatility in handling a variety of classification scenarios, especially in fine-grained tasks like avian classification. Moreover, our work contributes to the broader research community by developing the Bird-11K dataset, which encompasses a diverse range of bird species and presents a valuable resource for further explorations in fine-grained classification and beyond.
7 Limitations
Text encoder may not fully comprehend the bird descriptors
Our text encoder, pre-trained on a broad image-text dataset, may not fully capture the intricate details specific to birds. Furthermore, CLIP text encoders trained by contrastive learning are known to suffer from the binding problem and do not understand some logical operators such as “and”, “or”, or negation. PEEB accuracy depends directly on the quality of the text encoder.
Dependency on image encoder for part visibility
The image encoder’s role in determining the visibility of bird parts in an image poses another limitation. Our model operates under the assumption that 12 parts are always visible in a bird image, requiring it to score these parts even when they might not be visually present. In an ideal scenario, the model should learn to assign the absence part a low score. This approach, which lacks direct supervision, relies heavily on unsupervised learning derived from class labels. Consequently, the limited dataset size of approximately 290K training images in Bird-11K may not sufficiently support robust unsupervised learning.
Hallucinations in GPT-4 descriptors
The accuracy of our model is directly impacted by the quality of GPT-4 descriptors. Our empirical analysis across 20 bird classes revealed that, on average, 45% of these descriptors do not accurately reflect the birds’ features (Sec. F.1). However, we observed that revising certain descriptors in the CUB dataset led to a significant improvement of +10 points in classification accuracy for those classes (Sec. F.2). This primitive observation suggests that PEEB can be further improved if trained with human-labeled descriptors.
Acknowledgement
We are grateful to Cornell’s Macaulay Library for providing us with 55K images categorized into 10,534 bird species for the large-scale pre-training. We thank Pooyan Rahmanzadehgervi and Giang Nguyen from Auburn University for their helpful feedback on the early results. We also thank the valuable support and feedback from Phat Nguyen. This work is supported by NaphCare Foundations, Adobe gifts, and NSF grants (1850117, 2145767). AN is supported by NaphCare Foundations, Adobe gifts, and NSF grants (1850117, 2145767).
Appendix
Appendix A Architecture details
A.1 Image encoder and text encoder
We employ the image encoder and text encoder from OWL-ViT. In order to maintain a general understanding of natural languages and avoid overfitting our training samples, we keep the text encoder frozen for all training and experiments. This setup allows our design to be flexible about the choice of text encoder, e.g., one can easily replace the text encoder without changing other architecture.
A.2 Linear projection (for part embedding selection)
The image embedding will be forwarded to a Linear Projection layer, which is composed of a learnable logit scale, a learnable logit shift, and an Exponential Linear Unit (ELU) activation function. These processed image embeddings are then have the same dimension as the text embeddings. We select a single image embedding for each text query. In this context, the text queries correspond to the component names of the target object, which includes twelve distinct parts. This selection is based on the cosine similarity between the projected image embeddings and the text embeddings. Finally, the chosen images embeddings (before projection) will be sent to the Part MLP for classification and Box MLP for box prediction (Fig. A1, Step 1).
A.3 Part MLP
We introduce Part MLP to enable part-based classification. It comprises a three-layer MLP with GELU activations (Hendrycks and Gimpel, 2016). Part MLP takes in the selected part embeddings (i.e. output of step 1 in Fig. A1) and outputs a vector of size for each part, where is the dimension of descriptor embeddings. Part MLP is trained to map the selected part embeddings to the same dimensional space with descriptor embeddings to compute final logits for classification.
A.4 Box MLP
The Box MLP retained from OWL-ViT consists of a three-layer MLP. It takes the visual embedding as input and generates a four-element vector corresponding to the center coordinates and size of a bounding box (e.g., [x, y, width, height]). It is important to note that the image embedding inputs of Box MLP and Part MLP layers are the same, as shown in Fig. A1, Step 2.
A.5 Visual part embedding selection
As shown in Fig. A1 step 1, 1c, the image embeddings are first projected by a Linear Projection layer and compute the dot product with the encoded part names. The image embeddings (before linear projection) are chosen as visual part embeddings by selecting the embedding that has the highest similarity scores with the corresponding part after the linear projection.
A.6 Descriptor embedding matching
To enhance the model’s flexibility, we do not use a linear layer for classification. Instead, we adopt a strategy similar to CLIP: we compute the similarity matrix of the projected visual embeddings (image embeddings after processing by the Part MLP) and the text embeddings. Then, we sum the corresponding similarities of each part in the class; the class with the highest score is considered the predicted class as shown in Fig. A1, step 2, 2d. This design enables our proposed method to perform arbitrary ways of classification.
A.7 Implementation details
A.8 Training hyperparameters
A.9 Computational budget and infrastructures
We use 8 Nvidia RTX A100 GPUs for our experiments. The pertaining approximate takes 24 hours on Bird-11K. The fine-tuning takes 2 to 4 hours with one single GPU.
A.10 Training objectives
As discussed in Sec. 4.3, we have three objectives during the two training stages: (a) During the pre-training stage one, we contrastively pre-train the model to maximize the similarity between related part-descriptor pairs while minimizing the unrelated pairs using symmetric cross-entropy (CE) loss (Radford et al., 2021); (b) In pre-training stage two, we try to get rid of teacher model by mimic the teacher’s choice of the boxes with symmetric cross-entropy loss; (c) In stage two, we simultaneously train PEEB to improve box prediction with DERT losses (Zheng et al., 2020).
Pre-training stage one: Symmetric cross-entropy loss for contrastive pre-training
We first define the embeddings derived from the image and text encoders:
where is the input image, and is output image embeddings. Here, is the feature dimension of the image encoder. The text embedding is given by
where represents the tesxt input, and . In this case, is the feature dimension of the text encoder. The image embedding is then transformed by Part MLP layer (Fig. A1, 1b) to align its dimensions with the text embedding. This transformation is denoted as
where . The similarity matrix between the image and text embeddings is computed as the dot product of and the transpose of , expressed as
where . The image logits () and text logits () are then defined as
and
Next, we define the symmetric cross-entropy loss for the multi-modal embeddings.
where is the label for image and is the label for text.
Pre-training stage two: Symmetric cross-entropy loss to get rid of teacher model
To get rid of the teacher model, we train PEEB to mimic the feature selection between image embedding and part name embeddings (as shown in Fig. A1, 1c). We first binary the teacher logits and consider it as the ground truth label. Then, apply the same symmetric cross-entropy loss as described in eq. 7 with two minor differences: (1) The text input is part names rather than descriptions. (2) The Part MLP is replaced by Linear Projection (Fig. A1, 2c).
Fine-tuning: DETR losses
DETR losses are designed to optimize the box detection performance. We employ partial losses in our training for box predictions. Specifically, we employ corner-to-corner distance loss and GIoU loss. For the selected embeddings, we predict the boxes with Box MLP (Fig. A1, 2b)
where is the image selected image embeddings from eq. 1, is the predicted bounding boxes. Let be the ground truth boxes. The corner-to-corner distance loss is defined as
The GIoU loss is defined in Algorithm 1, and the total box loss is defined as
Step 2: (a) Encode input texts (i.e. N sets of 12-part descriptors) with the same text encoder to get . (b) Feed the selected part embeddings to box MLP to localize parts (in center format). (c) Also feed the selected part embeddings to part MLP to get in the same dimensional space with (d) Compute dot product between and , then diagonal sum for each class and over logits to get predicted label .
| Model | Epoch | Batch size | LR | Weight decay | # in-batch classes | Early stop | Training set | ||
| Train | Val | Train | Val | ||||||
| Pre-training stage 1 | |||||||||
| PEEB | 32 | 32 | 50 | 0.01 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 10 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 10 | Bird-11K | |
| Pre-training stage 2 | |||||||||
| PEEB | 32 | 32 | 50 | 0.01 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 5 | Bird-11K | |
| Model | Fine-tune from | Epoch | Batch size | LR | Weight decay | Early stop | Training set |
| PEEB | PEEB | 30 | 32 | 0.001 | 5 | CUB | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB ZSL (2015) | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB-SCS | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB-SCE | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | NABirds-SCS | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | NABirds-SCE |
Appendix B Model and dataset notations
B.1 Dataset notations
Following the conventional setup of ZSL, we execute certain exclusions to make sure none of the test classes or descriptors are exposed during pre-training. That is, Bird-11K and Bird-11K exclude all CUB and NABirds classes, respectively. For GZSL, we exclude all test sets in CUB, NABirds, and iNaturalist, denoted as Bird-11K. We provide detailed statistics for the three pre-training sets in Table A3.
| Training set | Number of images | Number of classes | ||
| Train | Val | Train | Val | |
| Bird-11K | 234,693 | 29,234 | 10,740 | 9,746 |
| Bird-11K | 244,182 | 28,824 | 10,602 | 9,608 |
| Bird-11K | 216,588 | 27,996 | 10,326 | 9,332 |
B.2 Model notations
We adopt a strategy based on the datasets excluded during training to simplify our model naming convention. Specifically:
PEEB is pre-trained model using Bird-11K datset.
PEEB is pre-trained model using the Bird-11K dataset.
PEEB is pre-trained model using the Bird-11K dataset.
We named fine-tuned models after the pre-trained model and the fine-tuned training set. For example, PEEB is fine-tuned from PEEB, on CUB training set.
Appendix C Generating part-based descriptors
CUB annotations initially comprise 15 bird parts. However, distinctions between the left and right part are not essential to our method, we merge them into a single part (i.e., “left-wing” and “right-wing” are merged into “wings”) Hence, we distilled the original setup into 12 definitive parts: back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail, throat. To compile visual part-based descriptors for all bird species within Bird-11K, we prompted GPT-4 (OpenAI, 2023) with the following input template:
A bird has 12 parts: back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail and throat. Visually describe all parts of {class name} bird in a short phrase in bullet points using the format ‘part: short phrase’
Where {class name} is substituted for a given bird name.
The output is a set of twelve descriptors corresponding to twelve parts of the query species. e.g. The response for Cardinal is:
Appendix D Datasets
D.1 Bird-11K
We provide a brief statistic of Bird-11K in Table A4. Bird-11K is a diverse and long-tailed avian dataset that only includes bird images. The descriptors generated by GTP4 are in English and only describe the visual features of the corresponding class. We propose Bird-11K for academic research only.
| Dataset | Images | Species |
| CUB-200-2011 (Wah et al., 2011) | 12,000 | 200 |
| Indian Birds (Vaibhav Rokde, 2023) | 37,000 | 25 |
| NABirds v1 (Van Horn et al., 2015) | 48,000 | 400 |
| Birdsnap v7 (Berg et al., 2014) | 49,829 | 500 |
| iNaturalist 2021-birds (Van Horn et al., 2021) | 74,300 | 1,464 |
| ImageNet-birds (Deng et al., 2009) | 76,700 | 59 |
| BIRDS 525 (Piosenka, 2022) | 89,885 | 525 |
| Macaulay Library at the Cornell Lab of Ornithology* | 55,283 | 10,534 |
| Bird-11K (Raw Data) | 440,934 | 11,097 |
| Bird-11K (pre-training set) | 294,528 | 10,811 |
Data splits
We provide data splits and metadata, e.g., file names, image size, and bounding boxes, along with the instruction of Bird-11K construction in our repository. Note that the Bird-11K dataset is for pre-training purposes; it is important to execute exclusion based on the test set.
License and terms
CUB (Wah et al., 2011): The dataset can be freely used for academic and research purposes; commercial use is restricted.
Indian Birds (Vaibhav Rokde, 2023): CC0: Public Domain.
NABirds v1 (Van Horn et al., 2015): For non-commercial research purposes, other use is restricted 333See Terms of Use here for detail: .
Birdsnap v7 (Berg et al., 2014): The dataset creator provides no specific license or terms of use. We only use this dataset for academic research until more specific details can be obtained.
iNaturalist 2021-birds (Van Horn et al., 2021): CC0: Public Domain.
ImageNet-birds (Deng et al., 2009): BSD-3-Clause license.
BIRDS 525 (Piosenka, 2022): CC0: Public Domain
Cornell eBird: We used the following 55,384 recordings from the Macaulay Library at the Cornell Lab of Ornithology. The data is for academic and research purposes only, not publicly accessible unless requested. (Please refer to our Supplementary Material for the full list):
ML187387391, ML187387411, ML187387421, ML187387431, ML262407521, ML262407481, ML262407531, ML262407491, ML262407511, ML257194111 ML257194071, ML257194081, ML257194061, ML495670791, ML495670781, ML495670801, ML495670771, ML183436431, ML183436451, ML183436441 ML183436411, ML183436421, ML256545901, ML256545891, ML256545841, ML256545851, ML256545831, ML169637941, ML238083081, ML169637881 ML169637911, ML238083111, ML238083051, ML169637971, ML299670841, ML64989231, ML299670831, ML64989241, ML299670791, ML64989251 ML246866001, ML246865941, ML246866011, ML246865961, ML246865971, ML333411961, ML240835531, ML240835541, ML240835701, ML240835591 ML245260391, ML245260341, ML245260371, ML245260411, ML245260421, ML245260431, ML245260441, ML240866351, ML240866331, ML240866321 ML240866341, ML240866371, ML248318661, ML248318571, ML248318591, ML248318581, ML248318631, ML245204281, ML245204311, ML245204371 ML245204381, ML245204291, ML245603571, ML245603521, ML245603511, ML245603491, ML245603501, ML245603601, ML245257771, ML245257651 ML245257631, ML245257661, ML245257761, ML247221051, ML247221061, ML247221071, ML247221081, ML240365811, ML240365751, ML240365781 ML240365761, ML300579541, ML247298551, ML247298541, ML247298561, ML247298611, ML247298571, ML247298591, ML247298601, ML247298631…
D.2 Dog-140
To pre-train PEEB on dogs, we construct Dog-140 by combining dog images from ImageNet21K and Stanford Dogs. Specifically, we selected 189 dog classes from ImageNet21K, and based on Fédération Cynologique Internationale (FCI) (Fédération Cynologique Internationale (2023), FCI), we merged them with 120 classes from Stanford Dogs, ending up with 140 classes. After merging, Dog-140 has 206,076 images in total. We provide a class distribution analysis in Fig. A4, where we can find that Dog-140 is roughly class-balanced.
Data splits
Similar to Bird-11K, we provide data splits and metadata, e.g., file names, image size, and bounding boxes, along with the instruction of Dog-140 construction in our repository.
License and terms
Appendix E Analysis
E.1 PEEB outperforms M&V in CUB and NABirds in ZSL setting
To rigorously evaluate the ZSL capabilities of our pre-trained models, we introduce a stress test on the CUB and NABirds datasets. The crux of this test involves excluding all classes from the target dataset (CUB or NABirds) during the pre-training. The exclusion ensures that the model has no prior exposure to these classes. Subsequently, we measure the classification accuracy on the target dataset, comparing our results against benchmarks set by CLIP and M&V in the scientific name test. In this experiment, we consider the scientific name test a ZSL test for CLIP and use them as the baseline because the frequencies of scientific names are much lower than common ones.
Experiment
To conduct this test, we pre-train our model on Bird-11K and Bird-11K, which deliberately exclude images bearing the same class label as the target dataset. Specifically, we test on our pre-train model PEEB and PEEB (see Table A1 for details), respectively.
Results
The primary objective is to ascertain the superiority of our pre-trained model, PEEB, against benchmarks like CLIP and M&V. For CUB, our method reported a classification accuracy of 17.9%, contrasting the 5.95% and 7.66% achieved by CLIP and M&V, respectively, as shown in Table A5. The PEEB score, which is marginally higher (+10) than M&V, highlights the advantages of our method that utilizes component-based classification. On the NABirds, our method surpassed the CLIP and M&V by (+1) point. The performance disparity between CUB and NABirds can be attributed to two factors: the elevated complexity of the task (555-way classification for NABirds versus 200-way for CUB) and the marked reduction in training data. An auxiliary observation, detailed in Sec. E.3, indicates that our pre-trained model necessitates at least 250k images to achieve admirable classification accuracy on CUB, but we only have 210k images training images in Bird-11K (Table A3).
| Method | CLIP | M&V | PEEB (ours) |
| CUB | 5.95 | 7.66 | 17.90 |
| NABirds | 4.73 | 6.27 | 7.47 |
E.2 Performance measurement on different noisy levels
In our evaluations, as indicated in Table 2, we discerned a marked performance disparity between the iNaturalist dataset and others. Probing this further, we identified image noise as a principal contributor to these discrepancies.
Experiment
A qualitative assessment of the iNaturalist test images revealed a significantly higher noise level than CUB or NABirds. To systematically study this, we utilize the object detector OWL-ViT to measure the size of the bird within the images. We formulated two filtered test sets based on the detector’s output, categorizing them by the bird’s size, specifically, the detected bounding box. Images were filtered out if the bird’s size did not exceed predetermined thresholds (areas of or pixels). Larger birds naturally reduced other content by occupying more image space, thus serving as a proxy for reduced noise. All three test sets, including the original, were evaluated using our pre-trained model PEEB.
Results
The results presented in Table A6 reveal a clear trend: as the image noise level decreases, the classification accuracy consistently improves, with gains ranging from (+6 to +17) points across the various methods. Notably, cleaner images consistently yield better results. At each noise level, our method outperforms the alternatives. While our method exhibits an impressive (+17 points) accuracy boost on the cleanest test set, this substantial gain also indicates that our model is sensitive to image noise.
| Splits | CLIP | M&V | PEEB (ours) |
| Original | 16.36 | 17.57 | 25.74 |
| pixels | 20.18 | 21.66 | 35.32 |
| pixels | 22.88 | 24.90 | 42.55 |
E.3 Number of training images is the most critical factor towards classification accuracy
Bird-11K, as shown in Fig. 4(a), is a highly imbalanced dataset characterized by a large amount of long-tailed classes. We conduct a comprehensive study to discern how variations in the number of classes and images affect the classification accuracy of our pre-trained models. Predictably, the volume of training images occurred as the most influential factor. However, a noteworthy observation was that the abundance of long-tailed data enhanced the model’s accuracy by approximately +1.5 points.
Experiment
We curated eight training sets based on varying class counts: 200, 500, 1,000, 2,000, 4,000, 6,000, 8,000, and 10,740. For each set, we maximized the number of training images. It is important to note that a set with a lesser class count is inherently a subset of one with a higher count. For instance, the 500-class set is a subset of the 2,000-class set. For each split, we apply the same training strategy as in Sec. 4.3.1, and choose the checkpoint with the best validation accuracy. We consider the CUB test set as a generic testing benchmark for all variants.
Results
As illustrated in Figure Fig. 4(b), there is a pronounced correlation between the increase in the number of images and the corresponding surge in accuracy. For instance, an increment from 106K to 164K images led to a rise in classification accuracy from 30.05% to 43.11%. The accuracy appears to stabilize around 60% when the image count approaches 250K. This trend strongly suggests that the volume of training images is the most critical factor for the pre-trained model. We believe that the accuracy of the pre-trained model could be further enhanced if enough data is provided. Interestingly, a substantial amount of long-tailed data bolsters the model’s performance, evident from +1.5 points accuracy improvement when comparing models trained on 2,000 classes to those on 10,740 classes. Note that the additional classes in the latter set averaged merely 2.2 images per class.
E.4 Ablation study on the influence of parts utilized
In this ablation study, we aimed to measure the impact of varying the number of distinct “parts” (back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail, and throat) used in our model. We experiment with a range from a single part to all 12 identifiable parts. Interestingly, even with a solitary part, the model could make correct predictions, though there was an evident decline in performance, approximately -20 points.
Experiment
Our testing ground is the pre-trained model PEEB, evaluated against the CUB test set. We assessed the model’s prowess utilizing various subsets of parts: 1, 3, 5, 8, and all 12. These subsets were derived based on the frequency of visibility of the parts within the CUB dataset, enabling us to compare the model’s performance when relying on the most frequently visible parts versus the least. For comparison, we also conduct a similar experiment on M&V, where we only use 1, 3, 5, 8, and 12 descriptors (if possible).
Results
Relying solely on the most frequent part led to a decline in classification accuracy by around -20 points, registering at 45.44% (Table A7). In contrast, utilizing the least frequent part resulted in a sharper drop of around -27, with an accuracy of 37.02%. As the model was furnished with increasing parts, its accuracy improved incrementally. The data underscores that optimal performance, an accuracy of 64.33%, is attained when all 12 parts are included. For M&V, the accuracy keeps increasing homogeneously from 5 to 12 descriptors, hinting that accuracy may increase further by increasing the number of descriptors.
| Number of Parts (descriptors) | 1 | 3 | 5 | 8 | 12 |
| Accuracy (most frequent parts) | 45.44 | 56.48 | 59.89 | 61.32 | 64.33 |
| Accuracy (least frequent parts) | 37.02 | 55.51 | 60.04 | 61.13 | 64.33 |
| Accuracy of Menon and Vondrick (2023) | 51.93 | 52.87 | 52.83 | 53.33 | 53.92 |
E.5 Training is essential for PEEB’s classification efficacy
In this ablation study, we highlight the pivotal role of training in the performance of PEEB on bird classification tasks. We demonstrate that without adequate tuning, the results are indistinguishable from random chance.
Experiment
We conduct the experiment based on OWL-ViT. We retain all components as illustrated in Fig. A1, with one exception: we substitute the part MLP with the MLP layer present in the box prediction head of OWL-ViT because the proposed layers require training. The MLP layers in the box prediction head project the part embeddings to match the dimensionality of the text embeddings. Our focus is on assessing the classification accuracy of the untuned PEEB on two datasets: CUB and NABirds.
Results
Table A8 reveals the outcomes of our experiment. Without training, PEEB yields classification accuracies of 0.55% for CUB and 0.31% for NABirds, both of which are proximate to random chance (0.5% for CUB and 0.1% for NABirds). However, with training, the model’s performance dramatically transforms: 64.33% for CUB (an increase of +63.78 points) and 69.03% for NABirds (a leap of +68.72 points) for PEEB. These pronounced disparities underscore the vital role of training in PEEB.
| CUB | NABirds | |
| PEEB (no training) | 0.55 | 0.31 |
| PEEB pre-trained | 64.33 | 69.03 |
| PEEB finetuned | 86.73 | - |
E.6 Failure analysis
Since PEEB has two branches, box detection, and descriptor matching, we would like to find out, in the failure case, what is the main cause. i.e., is it because of the mismatch in the descriptor to the part embeddings? Or is it because the box detection is wrong? From our ablation study, it turns out that most errors come from the descriptor-part matching.
Experiment
We conduct the experiment with PEEB on CUB test set. Specifically, we measure the box detection accuracy based on the key point annotation in CUB dataset, i.e., We consider the box prediction as correct if the prediction includes the human-annotated key point. We report the box prediction error rate (in %) based on parts.
Results
As shown in Table A9, the average error rate difference between success and failure cases is merely 0.38. That is, in terms of box prediction, the accuracy is almost the same, disregarding the correctness of bird identification. It indicates that the prediction error is predominantly due to the mismatch between descriptors and part embeddings. We also noted that some parts, like Nape and Throat, have a very high average error rate, which may greatly increase the matching difficulties between descriptors and part embeddings.
| Body Part | Average | Back | Beak | Belly | Breast | Crown | Forehead | Eyes | Legs | Wings | Nape | Tail | Throat |
| Failure Cases | 16.52 | 23.38 | 3.28 | 8.06 | 15.96 | 7.41 | 24.72 | 7.29 | 5.63 | 3.36 | 64.79 | 7.25 | 27.07 |
| Success Cases | 16.14 | 23.03 | 2.96 | 7.44 | 18.64 | 7.13 | 21.53 | 3.93 | 6.85 | 2.68 | 68.66 | 6.40 | 24.38 |
| Difference | 0.38 | 0.35 | 0.33 | 0.62 | -2.68 | 0.28 | 3.19 | 3.36 | -1.22 | 0.68 | -3.87 | 0.85 | 2.68 |
E.7 Evaluation of predicted boxes from PEEB
Our proposed method primarily aims to facilitate part-based classification. While the core objective is not object detection, retaining the box prediction component is paramount for ensuring model explainability. This section delves into an evaluation of the box prediction performance of our method against the OWL-ViT model.
Experiment
Given our focus on part-based classification, we aimed to ascertain the quality of our model’s box predictions. To this end, we employed two metrics: mean Intersection over Union (IoU) and precision based on key points. We opted for mean IoU over the conventional mAP because: (1) Ground-truth boxes for bird parts are absent, and (2) our model is constrained to predict a single box per part, ensuring a recall of one. Thus, we treat OWL-ViT’s boxes as the ground truth and evaluate the box overlap through mean IoU. Furthermore, leveraging human-annotated key points for bird parts, we measure the precision of predicted boxes by determining if they contain the corresponding key points. We evaluate our fine-tuned models on their corresponding test sets. For instance, PEEB, fine-tuned based on the CUB split (Akata et al., 2015), is evaluated on the CUB test set.
Results
Our evaluation, as presented in Table A10, shows that PEEB’s box predictions do not match those of OWL-ViT. Specifically, on average, there is a -5 to -10 points reduction in mean IoU for CUB and NABirds datasets, respectively. The disparity is less distinct when examining precision based on human-annotated key points; our method records about -0.14 points lower precision for CUB and -3.17 points for NABirds compared to those for OWL-ViT. These observations reinforce that while PEEB’s box predictions might not rival these dedicated object detection models, they consistently highlight the same parts identified by such models as shown in Fig. A6. It is important to note that our approach utilized the same visual embeddings for both classification and box prediction tasks. This alignment emphasizes the part-based nature of our model’s predictions.
| Models | Mean IoU | |||
| (1) All | (2) w/ Keypoints | Precision | ||
| CUB | OWL-ViT | 100.00 | 100.00 | 83.83 |
| OWL-ViT | 44.41 | 49.65 | 83.53 | |
| PEEB (Average) | 35.98 | 40.14 | 83.39 | |
| PEEB | 37.45 | 41.79 | 81.55 | |
| PEEB | 35.11 | 39.14 | 82.72 | |
| PEEB | 35.77 | 39.96 | 84.89 | |
| PEEB | 35.58 | 39.67 | 84.38 | |
| NABirds | OWL-ViT | 100.00 | 100.00 | 85.01 |
| OWL-ViT | 40.14 | 47.63 | 83.89 | |
| PEEB (Average) | 36.47 | 42.01 | 80.72 | |
| PEEB | 36.45 | 42.03 | 80.09 | |
| PEEB | 36.49 | 41.99 | 81.34 | |
Appendix F Study on GPT-4 generated descriptors
F.1 Noise measurement in GPT-4 generated descriptors
In this section, we conduct an empirical analysis to quantify the noise in descriptors generated by GPT-4 for 20 different classes within the CUB dataset. To achieve this, we manually inspect each descriptor and tally the instances where at least one factual error is present. Our findings reveal that every one of the 20 classes contains descriptors with errors, and on average, 45% of the descriptors necessitate corrections. This substantial noise level underscores the need for further refinement in our work, particularly in text descriptors.
We observe a notably high error rate in descriptors on the back and wings, with approximately 60% of these containing inaccurate information (refer to Table A11). This could be attributed to the challenges in distinguishing between the back and wings, given that the back is typically positioned behind the wings, yet exhibits considerable variability in size and shape. Addressing all descriptor issues by revising all 11,000 fine-grained descriptors would demand a significant investment of time and resources, which is beyond the scope of the current work. As such, we identify this as an area for future research and development, aiming to enhance the quality of the Bird-11K dataset.
| Back | Beak | Belly | Breast | Crown | Forehead | Eyes | Legs | Wings | Nape | Tail | Throat | Average | |
| Error Rate | 60 | 30 | 50 | 40 | 50 | 55 | 50 | 20 | 60 | 50 | 35 | 40 | 45 |
F.2 Revising descriptors improves classification accuracy
As mentioned in the limitation section, the descriptors are generated from GPT-4 and therefore noisy and incorrect. Given that PEEB accepts open vocabulary inputs for classification, a natural way to improve classification accuracy is to improve the correctness of the descriptors.
Experiment
We first collect descriptors of 183 CUB classes from AllAboutBirds. We then prompt GPT-4 to revise our original descriptors by providing the collected descriptor. We revise the descriptors with the following prompt:
Given the following descriptors of {class name}: {AllAboutBirds descriptors}. Can you revise the incorrect items below (if any) of this bird, return them as a Python dictionary, and use the key as the part name for each item? If a partś descriptor is not specifically described or cannot be inferred from the definition, use your own knowledge. Otherwise, leave as is. Note: please use a double quotation mark for each item such that it works with JSON format.
{Original descriptors}
Where {class name} the placeholder for the class name, {AllAboutBirds descriptors} is the description collected from AllAboutBirds, {Original descriptors} is the descriptors we used for training.
Due to the errors in the descriptors we used to train PEEB, simply replacing the descriptors with their revised version does not lead to better performance. Because the incorrect descriptors in training change the meaning of some of the phrases. For example, the belly of Blue bunting is pure blue, but the descriptors from GPT-4 is soft, creamy white. In addition, the GTP-4 uses the exact same descriptor in the belly for other classes, e.g., Blue breasted quail, which should be cinnamon. Blue Fronted Flycatcher, which should be yellow. Training the same descriptors with different colors confuses the model, and the model will convey the phrase “creamy white” with a different meaning to humans. Therefore, simply changing the descriptors to their’ revised version will not work. We empirically inspect the descriptors that PEEB can correctly respond to and replace the class descriptors with the revised version. Specifically, we replace the descriptors of 17 classes in CUB and test the classification accuracy on PEEB.
Results
As shown in Table A12, the overall accuracy increase +0.8 points. The average improvement of the revised class is around +10.8, hitting that if we have correct descriptors of all classes, we may significantly improve the classification accuracy of the pre-trained model. However, correcting all 11k class descriptors is too expensive and out of the scope of this work. We leave it as a further direction of improving the part-based bird classification.
| Descriptors | Original | Partially Revised | Avg. Improvement |
| PEEB | 64.33 | 65.14 | 10.80 |
Appendix G Qualitative Inspections
G.1 Visual comparison of predicted boxes
We provide a visual comparison of the box prediction from OWL-ViT, OWL-ViT, and PEEB in Fig. A6. We find that despite the fact that our predicted boxes have lower mean IoU compared to OWL-ViT, they are visually similar to the boxes as OWL-ViT.
OriginalPEEBOWL-ViTOWL-ViT
G.2 Qualitative examples of using randomized descriptors
We visually compare M&V and PEEB based on their utilization of descriptors. (Figs. A7, A8 and A9). Specifically, we randomly swap the descriptors of the classes and then use these randomized descriptors as textual inputs to the tested models to see how they perform. We observe that the scores from M&V tend to cluster closely together. Surprisingly, M&V’s prediction remains unchanged despite the inaccurate descriptors. In contrast, PEEB, when presented with randomized descriptors, attempts to identify the best match grounded on the given descriptors.
Original DescriptorRandom nonsense Descriptor
Original descriptorRandom nonsense descriptor
Original descriptorRandom nonsense descriptor
G.3 Examples of PEEB explanations for birds
Figs. A10, A11 and A12 are examples of how PEEB makes classification based on the descriptors and how it can reject the predictions made by M&V. Since we aggregate all descriptors for the final decision, even if some of them are similar in two classes, our method can still differentiate them from other descriptors. For instance, in Fig. A10, while other descriptors are similar, PEEB can still reject chesnut-sided warbler thanks to the distinct features of forehead, throat and belly.
G.4 Examples of PEEB explanations for dogs
Figs. A13, A14, A15, A16, A17, A18 and A19 are examples of how PEEB makes classification based on the descriptors in Stanford Dogs dataset. We showed that our model works well on dogs, which indicates that our proposed method is transferable to other domains while maintaining nice explainability as in birds.