### v3

# PEEB: Part-based Image Classifiers with an Explainable and Editable Language Bottleneck

Abstract
CLIP-based classifiers rely on the prompt containing a {class name} that is known to the text encoder.
Therefore, they perform poorly with new classes or the classes whose names rarely appear on the Internet (e.g., scientific names of birds).
For fine-grained classification, we propose PEEB—an explainable and editable classifier to (1) express the class name into a set of text descriptors that describe the visual parts of the class; and (2) match the embeddings of the detected parts with their textual descriptors in each class to compute a logit score for classification.
In a zero-shot setting where the class names are unknown,
PEEB significantly outperforms CLIP, achieving a 10-fold increase in top-1 accuracy.
Compared to part-based classifiers, PEEB not only achieves state-of-the-art (SOTA) accuracy in the supervised-learning setting—88.80% and 92.20% accuracy on CUB-200 and Dogs-120
, respectively—but also the first to enable users to edit the text descriptors to form a new classifier without any re-training.
Compared to concept bottleneck models, PEEB is also the SOTA in both zero-shot and supervised learning settings.
PEEB: Part-based Image Classifiers with an
Explainable and Editable Language Bottleneck
Thang M. Pham††thanks: Equal contribution† thangpham@auburn.edu Peijie Chen*† pzc0018@auburn.edu Tin Nguyen*† ttn0011@auburn.edu
Seunghyun Yoon § syoon@adobe.com Trung Bui § bui@adobe.com Anh Totti Nguyen † anh.ng8@gmail.com
†Auburn University §Adobe Research
1 Introduction
Fine-grained classification (Wah et al., 2011; Van Horn et al., 2015) is a long-standing computer-vision challenge. Furthermore, it is also important to explain how SOTA classifiers make a decision, e.g., which bird traits make a model think a given bird is Painted Bunting? (Fig. 1)
First, many bird classifiers claim to be explainable (Chen et al., 2019; Donnelly et al., 2022) by comparing the input image with a set of learned, part prototypes (Fig. 1b) or natural-language concepts (Fig. 1a). Yet, such prototypes are feature vectors and therefore not editable by users. Textual concepts are often compared against entire image for classification and it is unknown what image details match a given descriptor Menon and Vondrick (2023); Yang et al. (2023). Third, most vision-language classifiers need the prompt to have a known {class name} (like a special code instead of an expressive, natural description) that matches the input image Roth et al. (2023). Fourth, most classifiers require either training-set images in a supervised-learning setting or demonstration images in a zero-shot setting (Xian et al., 2018; Zhu et al., 2018). This requirement is impractical when building a classifier for a novel species whose photos do not yet exist in the database.
To address the above four problems, we propose PEEB, a Part-based image classifier that is Explainable and Editable via a natural-language Bottleneck. PEEB classifies images by grounding the textual descriptor of object parts provided by humans or GPT-4 (no images needed) to detected parts in the image (Fig. 1c). While PEEB leverages CLIP’s encoders (Radford et al., 2021), it uses no class names (e.g., Indigo Bunting) in the prompt. In contrast, CLIP-based models (Radford et al., 2021; Pratt et al., 2023; Menon and Vondrick, 2023) rely so heavily on the known class names that their accuracy drops significantly when the names are removed or replaced by less-common ones (Sec. 5.1).
For birds , we first define the parts of interest for identifying a bird.
We take the 15 parts defined in CUB Wah et al. (2011) and reduce them to 12 by merging similar parts, e.g. left wing and right wing are merged into wings.
Using GPT-4 (OpenAI, 2023), we construct a textual descriptor
to describe each bird part of every species (see Appendix C).
Next, PEEB localizes the 12 bird parts in the image and computes their matching scores with corresponding text descriptors (Fig. 2).
The sum of the 12 dot products between the paired visual and textual part embeddings would be the unnormalized distance (logits) between the input image and every class for classification
(Fig. 3).
For dogs
, we use a similar procedure.
To our knowledge, all existing public bird-image datasets (listed in Table A4) are limited in size (less than 100K images per dataset) and in diversity (less than 1,500 species per dataset), impeding large-scale, vision-language, contrastive learning. Therefore, for our pre-training, we construct Bird-11K, an exceptionally large dataset of bird images, comprising 290,000 images spanning across 11,000 species—essentially, all known bird species on Earth (Sec. 3). Bird-11K is constructed from seven existing bird datasets and 55,000 new images that we collect from the Macaulay Library. Similarly, we build Dog-140, a large-scale dataset of 206K dog images. Our main findings are:111Code & data: https://github.com/anguyen8/peeb
- 1.
2.
Our pre-trained PEEB outperforms CLIP-based classifiers by +8 to +29 percentage points (pp) in bird classification across CUB-200, NABirds-555, and iNaturalist-1486 (Sec. 5.2).
- 3.
4.
Compared with explainable CUB classifiers, PEEB scores an 88.80% top-1 accuracy, on par with the best CUB-200 classifiers (81–87% accuracy) that are trained via supervised learning and often not editable (Sec. 5.5).
5.
PEEB is applicable to multiple domains. On Stanford Dogs-120, PEEB scores 92.20%, substantially outperforming explainable models and on-par with SOTA black-box models (Sec. 5.6).
2 Related Work
Ante- vs. post-hoc explanations It is common to build fine-grained classifiers based on CNNs (He et al., 2016) or ViTs (He et al., 2022a). Although high-performing, these models do not admit an ante-hoc explanation interface (Gunning et al., 2021) and therefore rely on post-hoc interpretability methods, which tend to offer inaccurate and unstable, after-the-fact explanations (Rudin, 2019; Bansal et al., 2020). PEEB’s textual part-descriptors form an ante-hoc, natural-language explanation bottleneck that enables users to observe and edit the object attributes that contribute to each final prediction. By editing text descriptors, users can re-program the model without any further re-training (Fig. 2).
Prototypical Part Networks Like the explainable classifiers that learn part prototypes (Nauta et al., 2021; Donnelly et al., 2022; Nauta et al., 2022; Chen et al., 2019), PEEB also operates at the object-part level. However, there are two major differences. First, the textual part descriptors in PEEB are human-understandable and editable. In contrast, a part prototype Chen et al. (2019) is not directly editable or interpretable to users, and often interpreted by showing the nearest training-set image patches. Second, PEEB predicts a contextualized embedding for each object part and its spatial information can be viewed by inputting to the Box MLP (see Fig. 3) for bounding-box visualization.
Text-based Concept Bottlenecks Like PEEB, (Chen et al., 2020; Zhu et al., 2018; Rao et al., 2023; Paz-Argaman et al., 2020) also match visual part embeddings to text embeddings. Yet, they (1) do not use CLIP and instead rely on TF-IDF text features; (2) require a trained bird-part detector to detect 7 bird parts. In contrast, PEEB relies on CLIP, which admits easy text editability, and OWL-ViT, which serves as an open-vocabulary object-part detector that generalizes to many domains.
Recent vision-language models (VLMs) claim to be interpretable as they use textual concepts in the prompt. Yet, some works that rely on class-wise differential captions Esfandiarpoor and Bach (2023) or learned concept weights Yang et al. (2023); Panousis et al. (2023); Oikarinen et al. (2023); Yuksekgonul et al. (2023) do not generalize to unseen classes. The most recent, similar work to PEEB might be LaBo Yang et al. (2023), which; however, operates at the image level instead of patch level, and does not generalize to unseen classes.
Many CLIP-based classifiers (Han et al., 2023b; Pratt et al., 2023; Menon and Vondrick, 2023) rely heavily on having seen class names in the prompt and thus are neither explainable nor editable to users. Unlike CLIP-based models, PEEB reveals what image details are being used for classification by matching descriptors to corresponding visual object parts (e.g. a bird’s beak in Fig. 3).
Attribute-based Classifiers Attribute-Label Embedding (ALE) approaches (Akata et al., 2015; Yuksekgonul et al., 2023) employ a fixed set of attributes and train an attribute-to-label weight matrix for zero-shot classification. Several studies (Samuel et al., 2021; Xu et al., 2020; Hanouti and Le Borgne, 2023) highlight its effectiveness on datasets like CUB, SUN (Xiao et al., 2010), and AWA (Xian et al., 2019). Yet, in practice, ALE requires tabular data annotations for every new class in the dataset (e.g., 312 attributes per CUB species), editing the weight matrix, and model re-training. In contrast, to add an unseen class to PEEB, users would only need to describe its 12 bird parts in natural language.
3 Datasets
3.1 Test classification benchmarks
3.2 Bird-11K dataset construction
We combine labeled images from 7 distinct datasets and an extra 55K images (10,534 classes) from Cornell’s Macaulay Library, to form a unified Bird-11K dataset222We do not redistribute the published datasets but release a script for reconstructing Bird-11K on Github. (Sec. D.1) for large-scale pre-training. To the best of our knowledge, Bird-11K, comprising 440,934 images spanning 11,183 classes, is the first bird dataset to encompass almost all species on Earth. Since PEEB learns to match visual parts with textual descriptors, it requires that bird images be distinctly visible and sufficiently large for accurate part localization and matching (see Sec. E.3 for ablation studies). However, small and “hard-to-see” bird images in Bird-11K make the dataset noisy and the training complex. Thus, we harness OWL-ViT (Minderer et al., 2022) to detect a bird in all images using the prompt “bird” and filter out images where the detected bird’s bounding box is smaller than pixels. We find OWL-ViT’s bird detections to be fairly accurate—its mean Intersection over Union (IoU) between the predicted bird boxes and ground-truth boxes on CUB dataset is 0.91.
As class labels from different sources are either general (e.g. Cardinal) or fine-grained (e.g. Yellow vs. Northern Cardinal), we retain only the fine-grained species for more diverse training and exclude all general classes to avoid label ambiguity. Following these filtering steps, the refined Bird-11K dataset retains 294,528 images across 10,811 classes (Table A4).
For each species in Bird-11K, we generate a set of part-based descriptors using GPT-4 (Appendix C). These generated descriptors (see Fig. 4) may not be 100% accurate but discriminative enough to help GPT-4V reach 69.40% accuracy on the CUB-200 test set (Table 3). That is, in the same prompt, we feed each test image along with the 200 CUB classes’ part-based descriptors and ask GPT-4V to select a matching class label for (details in Sec. F.1).
3.3 Dataset splits for contrastive pre-training
There are two common settings in the zero-shot learning literature—standard zero-shot (ZSL) and generalized zero-shot (GZSL).
ZSL is a stricter setup where a model is only tested on the classes unseen during any prior training. We ensure test-set classes from datasets (e.g., CUB-200 or NABirds-555) are not seen during pre-training. For example, to test on CUB under ZSL, we exclude all 200 CUB classes and their images from our pre-training on Bird-11K.
Following the ZSL literature, we use the CUB split proposed by Akata et al. (2015) and two harder splits: Super-Category-Shared/Exclusive (SCS/SCE) by Elhoseiny et al. (2017). For example, in ZSL on CUB, we exclude all CUB classes in Bird-11K for pre-training and finetune only on the corresponding training set given by a ZSL split.
GZSL is closer to the real-world setup where models are tested on both seen & unseen classes. CLIP’s “zero-shot” tests technically fall under GZSL as its Internet-scale training set might actually have images from the test classes. To test PEEB under GZSL, we exclude the test sets of CUB, NABirds, and iNaturalist, and directly evaluate the Bird-11K-pretrained models without further finetuning.
4 Method
4.1 Backbone: OWL-ViT object-part detector
OWL-ViT is an open-vocabulary detector that detects objects and parts in an image given a text prompt, even if the model is not explicitly finetuned to detect those concepts. OWL-ViT consists of four networks (Fig. 3): (1) a ViT-based image encoder, (2) an architecturally identical text encoder, (3) a bounding-box regression head called Box MLP, and (4) and a Linear Projection. Box MLP is a three-layer Multilayer Perceptron (MLP) with GELU activations (Hendrycks and Gimpel, 2016) after each of the first two layers. Linear Projection maps the visual and text embeddings to the same space (see Fig. 1 in Minderer et al. (2022)).
4.2 PEEB classifier
Architecture PEEB (Fig. 3) has five networks: an image encoder, a text encoder, a Linear Projection, a Part MLP, and a Box MLP.
We introduce Part MLP to map the visual and textual part embeddings to the same space for computing dot products (logits) for classification ( in Fig. 3). This design allows PEEB to easily extend the number of classes without any re-training. Except for Part MLP, all components are adopted from the OWL-ViT framework. Details of all components are in Appendix A.
Inference Given an input image, we first use the 12 generic part names to select the visual part embeddings based on cosine similarity. These selected visual part embeddings are then simultaneously fed into both Part MLP and Box MLP.
Box MLP predicts the bounding box from each part embedding. We compute a dot product to measure the similarity between each embedding output from Part MLP and a corresponding part-descriptor embedding. For classification, a class logit is the sum of the 12 dot products, which essentially computes the similarity between the 12 parts in the image and the 12 text descriptors of each class.
4.3 Training strategy
Trainable networks In preliminary experiments, we find training only Part MLP (while keeping all other networks frozen) to result in poor accuracy. Therefore, we train Part MLP from scratch and also finetune the image encoder, Linear Projection, and Box MLP. We finetune all OWL-ViT components from their original weights. In contrast, our proposed Part MLP starts from random weights. Our training has two phases: (a) 2-stage pre-training on the large-scale Bird-11K dataset and (b) finetuning on downstream tasks. More hyperparameter details are in Sec. A.8.
Objectives We aim to train PEEB to classify images well while maintaining the ability to detect object parts. This translates into three training objectives: (1) Train the Part MLP contrastively using a symmetric cross-entropy (SCE) loss (Radford et al., 2021) to maximize the similarity between region-text pairs while minimizing the similarity for negative pairs; (2) Train the Linear Projection using a SCE loss to mimic OWL-ViT’s behaviors (i.e. the similarity matrix) for part selection; and (3) Train Box MLP to predict bounding boxes with DETR losses (Zheng et al., 2021) i.e. a linear combination of corner-to-corner distance loss and GIoU loss (Rezatofighi et al., 2019).
All three losses are described in Sec. A.10.
A challenge when jointly minimizing all three losses above is that PEEB’s validation loss improves significantly slowly perhaps because of some tension between the two SCE losses and the DETR detection loss. To overcome this challenge, we split the pre-training phase into two stages: (1) first, train the image encoder and Part MLP for classification using the SCE loss; then (2) train the Linear Projection and Box MLP using the 2nd and 3rd loss so they can adapt their weights to the updated image encoder. We always keep the text encoder frozen since we want to preserve its generalizability to the descriptors of unseen objects.
4.3.1 2-stage pre-training on Bird-11K
Stage 1: Contrastive learning The image encoder and Part MLP are jointly trained using a SCE loss, which allows PEEB to learn to map the visual parts to corresponding text descriptors. In this stage, we use a pre-trained OWL-ViT to select 12 part embeddings per input image (i.e., teacher forcing) to ensure the selection of part embeddings is meaningful and consistent while the embeddings themselves are updating (see Fig. A2).
Stage 2: Learning to detect from a teacher After the image encoder is modified in Stage 1, we then train the Linear Projection and Box MLP jointly. We use the OWL-ViT as the teacher to train both components. Using SCE loss, we train the Linear Projection such that the similarity matrix between the part-names and visual parts matches those of the teacher (Fig. A3, 1a–c, 2a–c). Given the absence of human-annotated boxes for object parts, we train Box MLP to predict the same boxes as the predictions by OWL-ViT using DETR losses (Fig. A3, 2d). In this Stage 2, the image encoder is frozen while Part MLP is not involved.
After 2-stage training, PEEB can perform zero-shot classification and generate explanations.
4.3.2 Finetuning on classification tasks
We finetune the pre-trained PEEB on downstream tasks (CUB, NABirds and iNaturalist for birds and Dogs-120 for dogs) to further improve its accuracy. In this phase, to adapt to a downstream task, all components except the text encoder are trained jointly and the loss for Part MLP is changed from SCE (contrastive) to CE (classification) while the other two losses (DETR) are kept intact.
5 Experiments & Results
We conduct extensive experiments to evaluate PEEB on multiple bird datasets (CUB, NABirds, iNaturalist) and on GZSL (Secs. 5.1 and 5.3), ZSL (Sec. 5.4) and also supervised learning settings.
We also find PEEB to perform well on
dog image classification on Dogs-120 (Sec. 5.6).
Original Descriptors (a)Randomized, wrong descriptors (b)
Blue JayBlue Jay 0.0059Blue Jay 0.0058
M&V Menon and Vondrick (2023)
Blue Jay 0.6899 (c)Least Tern 0.0611 (d)
PEEB (ours)
5.1 CLIP-based classifiers rely mostly on {class names} (not descriptors)
M&V show that inserting extra GPT-3-generated descriptors into CLIP’s prompts increases its accuracy on many tasks Menon and Vondrick (2023). Yet, it is unknown how important these expressive descriptors are compared to the class names. To answer this question, we conduct two experiments on all three models: CLIP, M&V, and our PEEB.
Experiment 1 We evaluate the role of expressive descriptors to CLIP-based models and to PEEB by measuring the drop in CUB-200 accuracy of each model when the descriptors are randomized.
For M&V and PEEB, we randomize the descriptors by swapping each descriptor with another from an arbitrary class (examples in Fig. 4).
Experiment 2 We test the dependence of models on class names by measuring the accuracy drop when they are replaced by scientific names (e.g., Painted Bunting Passerina ciris) on CUB, NABirds, and iNaturalist.
Results When random descriptors are used, M&V accuracy drops marginally by -0.9 pp (Table 1; 53.70% 52.88%), showing that descriptors actually play a minimal role in model predictions. Instead, CLIP and M&V mostly rely on class names (e.g., 53.78% 7.66%; Table 2)—their accuracy drops drastically when class names are replaced by scientific names, which are less common.
In contrast, the expressive part descriptors play a major role in PEEB whose accuracy decreases significantly to near random-chance (64.33% 0.88%; Table 1) when the descriptors are randomized. Indeed, in PEEB, the textual descriptors serve as editable and interpretable model parameters that can be refined and extended by humans to account for new classes (Fig. 2).
5.2 Pre-trained PEEB outperforms CLIP-based classifiers in GZSL
The dependence on class names (Sec. 5.1) suggests that CLIP was exposed to these names during training. Thus, for a fair comparison, we compare PEEB with CLIP-based classifiers in the GZSL setting.
Experiment We train PEEB on Bird-11K using the two-stage pre-training (described in Sec. 4.3.1), and then test it on CUB, NABirds, and iNaturalist without any finetuning. That is, PEEB’s contrastive pre-training is at the part level and therefore the model has not seen the species labels of images.
Results PEEB outperforms both CLIP and M&V on all three datasets by huge margins of around +10, +28, and +8 pp on CUB-200, NABirds-555 and iNaturalist-1486, respectively (see Table 2).
5.3 PEEB is superior to text descriptor-based classifiers in GZSL on CUB-200
The advent of CLIP (2021) by OpenAI enabled a class of image classifiers that match the input image with pre-defined textual prompts that may include class names or descriptors of the classes. Yet, in contrast to PEEB, these descriptors often describe the entire image and are also matched (via dot product) with the entire image instead of image regions. Here, we compare PEEB with these methods in the GZSL setting on CUB-200.
Experiment We repeat the same experiments in Sec. 5.2. As these bird classifiers (listed in Table 3) were reported on CUB only (not NABirds or iNaturalist), our comparison is on CUB.
Results PEEB exhibits superior GZSL performance, outperforming recent text concept-based approaches by +3 to +10 pp (Table 3b). Compared to prior methods, PEEB is the only one to detect visual object parts and match them with text descriptors. Furthermore, attribute-based classifiers, e.g., Yuksekgonul et al. (2023) require re-training to adapt to new classes or datasets (e.g., NABirds or iNaturalist) in the same domain. In contrast, to apply PEEB to NABirds or a new class, no training is required—it is necessary to only edit its text descriptors (see Fig. 2). Interestingly, PEEB is 2nd-best model, only after GPT-4V (64.33% vs. 69.40%), which is given the same textual part descriptors as PEEB for all 200 CUB classes and asked to select a matching class for each image.
* 1-shot learning. † -means with .
| Method | Acc (%) | {c} | Textual descriptors |
| (a) Vision-language models with class names {c} in the prompt | |||
| CLIP (2021) | 52.02 | ✓ | Image-level |
| M&V (2023) | 53.78 | ✓ | Image-level |
| FuDD (2023) | 54.30 | ✓ | Image-level |
| Han et al. (2023b) | 56.13 | ✓ | Image-level |
| (b) Vision-language models with text bottlenecks and no class names {c} | |||
| LaBo (2023) | 54.19† | ✗ | Image-level |
| Yan et al. (2023) | 60.27* | ✗ | Image-level, attribute-based |
| PEEB (ours) | 64.33 | ✗ | Part-level |
| GPT-4V (2023) | 69.40 | ✓ | Part-level |
| (c) Concept-Bottleneck Models with attribute-based, non-textual bottlenecks | |||
| CBM (2020) | 62.90 | ✗ | Attribute-based, tabular data |
| PCBM (2023) | 61.00 | ✗ | Attribute-based, tabular data |
5.4 PEEB generalizes to traditional ZSL
Since PEEB outperforms modern vision-language models in GZSL (Sec. 5.3), we are motivated to further compare PEEB with SOTA approaches in the traditional ZSL setting (where the test classes are excluded from all prior training).
Experiment We evaluate PEEB on two common ZSL splits: (a) the CUB split Akata et al. (2015); and (b) the Super-Category-Similar/Exclusive (SCS/SCE) splits Elhoseiny et al. (2017) on CUB and NABirds. The SCS (Easy) and SCE (Hard) splits are designed to test two generalization levels (generalizing to close vs. distant unseen species).
Aligned with ZSL conventions, we exclude all species that exist in CUB or NABirds from the pre-training and then finetune PEEB using the train/test splits by Akata et al. and Elhoseiny et al.. We randomly take 10% of the training set as the validation set and choose the checkpoints based on the lowest validation loss.
| Methods | CUB | NABirds | ||||
| Seen | Unseen | Mean | Seen | Unseen | Mean | |
| (a) Data split by Akata et al. (2015) | ||||||
| CLORECLIP (2023a) | 65.80 | 39.10 | 49.05 | n/a | ||
| PEEB (ours) | 80.78 | 41.74 | 55.04 | |||
| (b) SCS/SCE splits by Elhoseiny et al. (2017) | ||||||
| SCS | SCE | Mean | SCS | SCE | Mean | |
| (Easy) | (Hard) | (Easy) | (Hard) | |||
| S2GA-DET (2018) | 42.90 | 10.90 | 17.38 | 39.40 | 9.70 | 15.56 |
| GRZSL (2018) | 44.08 | 14.46 | 21.77 | 36.36 | 9.04 | 14.48 |
| ZEST (2020) | 48.57 | 15.26 | 23.22 | 38.51 | 10.23 | 16.17 |
| CANZSL (2020) | 45.80 | 14.30 | 21.12 | 38.10 | 8.90 | 14.43 |
| DGRZSL (2021) | 45.48 | 14.29 | 21.75 | 37.62 | 8.91 | 14.41 |
| DPZSL (2023) | 45.40 | 15.50 | 23.11 | 40.80 | 8.20 | 13.66 |
| PEEB (ours) | 44.66 | 20.31 | 27.92 | 28.26 | 24.34 | 26.15 |
Results By a large margin, PEEB outperforms CLORECLIP, a SOTA CUB method in the (2015) split, on both seen and unseen classes (Table 4a). On the (2017) splits, PEEB is the SOTA in the Hard set on both CUB and NABirds datasets (Table 4b). That is, PEEB is better in generalizing to distant, unseen classes. This may be because PEEB decomposes both the image and the text descriptors into part-level features, which can re-combine to match an arbitrary unseen class (as illustrated in Fig. 2).
Interestingly, on both CUB and NABirds, PEEB is competitive but not SOTA on the Easy sets (Table 4b; Easy)—those classes that are close to the training-set classes and thus considered easier to identify. Overall, considering the harmonic mean over both Easy and Hard accuracy scores, PEEB is the SOTA on both CUB and NABirds.
5.5 Finetuning the pre-trained PEEB on CUB-200 yields a competitive explainable classifier in supervised learning
After finding that PEEB performs well in both GZSL (Sec. 5.3) and ZSL settings (Sec. 5.4), here we test finetuning the pre-trained PEEB on CUB-200. That is, we compare PEEB against SOTA explainable classifiers in the supervised learning setting to gain insights into our method’s adaptability to downstream tasks.
Experiment To understand the impact of pre-training and image resolution, we test finetuning three different PEEB variants: (1) PEEB initialized from OWL-ViT without pre-training on Bird-11K; (2) PEEB initialized from OWL-ViT with pre-training (described in Sec. 5.2); and (3) PEEB initialized from OWL-ViT with pre-training. We take each PEEB model and finetune all components on CUB-200, for 30 epochs with a batch size of 30, a learning rate of . Detailed hyperparameters are in Table A2.
Results Without pre-training, PEEB reaches 77.80% top-1 accuracy on CUB-200. Yet, first pre-training on Bird-11K and then finetuning on CUB yields 86.73%, the best among all explainable classifiers (Table 5b–c). Besides, pre-training PEEB from the higher-resolution OWL-ViT results in a further gain of +2.07 (86.73% 88.80%), which is intuitive since fine-grained classification is known to benefit from higher resolutions.
For a complete assessment, we compare and find PEEB to underperform SOTA standard, black-box classifiers by a few points (Table 5a).
| Methods | Model size | Backbone | Acc (%) |
| (a) SOTA black-box classifiers | |||
| Base (ViT) (2021) | 22M | DeiT-S (2021) | 84.28 |
| ViT-Net (2022a) | 26M | DeiT-S | 90.10 |
| (b) Concept-bottleneck classifiers | |||
| CBM Koh et al. (2020) | 11M | ResNet-18 | 80.10 |
| CPM Panousis et al. (2023) | 155M | ViT-B/16 | 72.00 |
| CDM Oikarinen et al. (2023) | 155M | ViT-B/16 | 74.31 |
| LaBo Yang et al. (2023) | 427M | ViT-L/14 | 81.90 |
| (c) Part-based, explainable classifiers | |||
| ProtoPNet (2019) | 22M | DeiT-S | 84.04 |
| ProtoTree (2021) | 92M | ResNet-50 | 82.20 |
| TesNet (2021) | 79M | DenseNet-121 | 84.80 |
| Deformable ProtoPNet (2022) | 23M | ResNet-50 | 86.40 |
| ProtoPFormer (2022) | 22M | DeiT-S | 84.85 |
| PEEB (ours) | 155M | ||
| pre-training + finetuning only | 155M | OWL-ViT | 77.80 |
| pre-training + finetuning | 155M | OWL-ViT | 86.73 |
| pre-training + finetuning | 155M | OWL-ViT | 88.80 |
5.6 Applying PEEB to
dog identification
We have found that our pre-training dataset construction and PEEB performs well for bird identification. By design, our method is not specific to birds but is instead applicable to any fine-grained classification domains assuming that the object is decomposable into parts. Here, we show that our method performs well on dog image classification as well.
Pre-training dataset construction First, we define a set of six dog parts that humans use to identify dog species. We use all 4 dog parts defined by PartImageNet He et al. (2022b)—head, body, legs, and tail—and two more parts—muzzle and ears—based on our manual image examination.
We combine ImageNet-21K and Stanford Dogs-120 into Dog-140, our large-scale pre-training dataset spanning 140 dog species (details in Sec. D.2). For each class, we prompt GPT-4 to get the descriptors for 6 parts. For each image in Dog-140, we run OWL-ViT to detect the corresponding boxes for 6 pre-defined parts.
Experiment Following the supervised learning experiment in Sec. 5.5, we first we pre-train PEEB (initialized from OWL-ViT) on Dog-140 and then further finetune it on Dogs-120.
Results Finetuning PEEB on Dogs-120 from OWL-ViT without pre-training on Dog-140 results in a 74.17% top-1 accuracy on Dogs-120 (Table 6b). In contrast, pre-training on Dog-140 only without finetuning results in much better Dogs-120 accuracy of 87.38%. That is, our contrastive pre-training helps model generalize (in a GZSL setting) while directly finetuning on Dogs-120 perhaps yields an overfitting model. Yet, pre-training and then finetuning reaches the best supervised learning accuracy of 92.20%, which is SOTA among all explainable models reported on Dogs-120.
Besides, PEEB offers novel, editable image-text grounding explanations (see Fig. 5).
| Methods | Model size | Backbone | Acc (%) |
| (a) SOTA black-box classifiers | |||
| TransFG (2022a) | 86M | ViT-B/16 | 92.30 |
| ViT-Net (2022b) | 86M | DeiT-B | 93.60 |
| SR-GNN (2022) | 32M | Xception | 97.00 |
| (b) Explainable methods | |||
| FCAN (2016) | 50M | ResNet-50 | 84.20 |
| RA-CNN (2017) | 144M | VGG-19 | 87.30 |
| ProtoPNet (2019) | 22M | DeiT-S | 77.30 |
| Deformable ProtoPNet (2022) | 23M | ResNet-50 | 86.50 |
| PEEB (ours) | 155M | ||
| pre-training + finetuning only | 155M | OWL-ViT | 74.17 |
| pre-training + finetuning | 155M | OWL-ViT | 87.37 |
| pre-training + finetuning | 155M | OWL-ViT | 92.20 |
6 Discussion and Conclusion
We introduce PEEB, a unique, novel explainable classifier due to its editability (Fig. 2) and operation at the part level on both image and text sides. The part-level operation makes PEEB applicable to fine-grained classification. Yet, it is also interesting to extend PEEB into an object-level model for multi-domain tasks like ImageNet or VQA.
Besides enabling users to edit PEEB’s text descriptors to re-program PEEB, it might also be promising to let users edit the bounding boxes while working with PEEB to improve the human-AI team accuracy Nguyen et al. (2024). On object detection, PEEB’s Box MLP performs on-par with OWL-ViT based on quantitative (Sec. E.7) and qualitative results (Appendix G).
Finally, we contribute to the broader research community by curating the Bird-11K and Dog-140 datasets and showing that it is possible to leverage them for large-scale training.
7 Limitations
Text encoder may not fully comprehend the bird descriptors
Our CLIP text encoder, pre-trained on an Internet-scale image-text dataset Radford et al. (2021), may not fully capture the intricate details specific to birds. Furthermore, the CLIP text encoder is known to suffer from the binding problem and do not understand some logical operators such as “and”, “or”, or negation. PEEB accuracy depends heavily on the quality of the text encoder.
Assumption that object parts mostly visible
PEEB operates based on the assumption that most if not all of the object parts are visible in the image. In cases where a part is missing or occluded, the model may still assign a non-zero similarity score (i.e. a non-zero dot product between the image-part embedding and its associated text descriptor), which makes it harder to separate classes. It might be beneficial to incorporate extra training samples and specifically encourages PEEB to assign zero image-text similarity score to the missing or occluded parts.
Hallucinations in GPT-4 descriptors
The accuracy of PEEB is directly governed by the accuracy of descriptors, which are currently generated by GPT-4. Yet, our manual assessment over 20 bird classes reveals that, on average, 45% of these descriptors do not accurately reflect the birds’ features (Sec. F.2). Also, we observe that revising certain descriptors in the CUB dataset led to a significant improvement of +10 points in classification accuracy for those classes (Sec. F.3). This primitive observation suggests that PEEB can be further improved if trained with more accurate, human-labeled descriptors.
Acknowledgement
We are grateful to Cornell’s Macaulay Library for providing us with 55K images categorized into 10,534 bird species for the large-scale pre-training. We thank Pooyan Rahmanzadehgervi, Giang Nguyen, Tin Nguyen, and Hung Huy Nguyen from Auburn University for their helpful feedback on our early results. We also thank Phat Nguyen for her valuable support and feedback. AN is supported by NaphCare Foundations, Adobe gifts, and NSF grant no. 2145767.
Appendix for:
PEEB: Part-based Image Classifiers with an
Explainable and Editable Language Bottleneck
Appendix A Architecture details
A.1 Image encoder and text encoder
We employ the image encoder and text encoder from OWL-ViT. In order to maintain a general understanding of natural languages and avoid overfitting our training samples, we keep the text encoder frozen for all training and experiments. This setup allows our design to be flexible about the choice of text encoder, e.g., one can easily replace the text encoder without changing other architecture.
A.2 Linear projection (for part embedding selection)
The image embedding will be forwarded to a Linear Projection layer (see detail implementation here), which is composed of a learnable logit scale, a learnable logit shift, and an Exponential Linear Unit (ELU) activation function. These processed image embeddings then have the same dimension as the text embeddings. For OWL-ViT, the image embeddings are projected from 768 to 512. We select a single image embedding for each text query. In this context, the text queries correspond to the component names of the target object, which includes twelve distinct parts. This selection is based on the cosine similarity between the projected image embeddings and the text embeddings. Finally, the chosen images embeddings (before projection) will be sent to the Part MLP for classification and Box MLP for box prediction (Fig. A1, Step 1).
A.3 Part MLP
We introduce Part MLP to enable part-based classification (see implementation detail here). It comprises a three-layer MLP with GELU activations (Hendrycks and Gimpel, 2016) . Part MLP takes in the selected part embeddings (i.e. output of step 1 in Fig. A1) and outputs a vector of size for each part, where is the dimension of descriptor embeddings (for OWL-ViT, the input dimension is 768, and ). Part MLP is trained to map the selected part embeddings to the same dimensional space with descriptor embeddings to compute final logits for classification.
A.4 Box MLP
The Box MLP retained from OWL-ViT consists of a three-layer MLP (see here for implementation detail). It takes the visual embedding as input and generates a four-element vector corresponding to the center coordinates and size of a bounding box (e.g., [x, y, width, height]). It is important to note that the image embedding inputs of Box MLP and Part MLP layers are the same, as shown in Fig. A1, Step 2.
A.5 Visual part embedding selection
As shown in Fig. A1 step 1, 1c, the image embeddings are first projected by a Linear Projection layer and compute the dot product with the encoded part names. The image embeddings (before Linear Projection) are chosen as visual part embeddings by selecting the embedding that has the highest similarity scores with the corresponding part after the Linear Projection.
Step 1: (a) Encode an input image and texts (i.e. 12 part names) by the image and text encoder to get patch embeddings and text embeddings . (b) Feed to Linear Projection to get in the same dimensional space with and compute dot product between and . (c) over embeddings to select 12 part embeddings.
Step 2: (a) Encode input texts (i.e. N sets of 12-part descriptors) with the same text encoder to get . (b) Feed the selected part embeddings to Box MLP to localize parts (in center format). (c) Also feed the selected part embeddings to Part MLP to get in the same dimensional space with (d) Compute a dot product between and , then diagonal sum for each class and over logits to get predicted label .
A.6 Descriptor embedding matching
To enhance the model’s flexibility, we do not use a linear layer for classification. Instead, we adopt a strategy similar to CLIP: we compute the similarity matrix of the projected visual embeddings (image embeddings after processing by the Part MLP) and the text embeddings. Then, we sum the corresponding similarities of each part in the class; the class with the highest score is considered the predicted class as shown in Fig. A1, step 2, 2d. This design enables our proposed method to perform arbitrary ways of classification.
A.7 Implementation details
A.8 Training hyperparameters
A.9 Computational budget and infrastructures
We use 8 Nvidia RTX A100 GPUs for our experiments. The pertaining approximate takes 24 hours on Bird-11K. The finetuning takes 2 to 4 hours with one single GPU.
A.10 Pre-training and Finetuning objectives
As discussed in Sec. 4.3, we have three objectives during the Pre-training phase:
- 1.
2.
Pre-training Stage 2: (Fig. A3) We try to remove the dependence on the OWL-ViT teacher model by training PEEB to mimic OWL-ViT’s box predictions using the SCE loss.
- 3.
During the Finetuning phase where we finetune on a downstream task (e.g. Dogs-120 or CUB-200), we also employ the same three losses. However, we change the first loss from SCE into CE since on the downstream classification task, the classifier is tasked with selecting one class that matches the single input image from a set of classes.
A.10.1 Pre-training stage one: Symmetric cross-entropy loss for contrastive pre-training
We first define the embeddings derived from the image and text encoders:
where is the input image, and is output image embeddings. Here, is the feature dimension of the image encoder. The text embedding is given by
where represents the tesxt input, and . In this case, is the feature dimension of the text encoder. The image embedding is then transformed by Part MLP layer (Fig. A1, 1b) to align its dimensions with the text embedding. This transformation is denoted as
where . The similarity matrix between the image and text embeddings is computed as the dot product of and the transpose of , expressed as
where . The image logits () and text logits () are then defined as
and
Next, we define the symmetric cross-entropy loss for the multi-modal embeddings.
where is the label for image and is the label for text.
A.10.2 Pre-training stage 2: Symmetric cross-entropy loss to mimic the teacher OWL-ViT detector
To mimic the object detection capability of the OWL-ViT teacher, we train PEEB to mimic the image-text similarity prediction between image embedding and textual part-name embeddings (as shown in Fig. A1, 1c). We first binary the teacher logits and consider it as the ground truth label. Then, apply the same symmetric cross-entropy loss as described in eq. 7 with two minor differences: (1) The text input is part names rather than descriptions. (2) The Part MLP is replaced by Linear Projection (Fig. A1, 2c).
A.10.3 Pre-training stage 2: DETR losses to mimic the teacher OWL-ViT detector
DETR losses are designed to optimize the box detection performance. We employ partial losses in our training for box predictions. Specifically, we employ corner-to-corner distance loss and GIoU loss. For the selected embeddings, we predict the boxes with Box MLP (Fig. A1, 2b)
where is the image selected image embeddings from eq. 1, is the predicted bounding boxes. Let be the ground truth boxes. The corner-to-corner distance loss is defined as
The GIoU loss is defined in Algorithm 1, and the total box loss is defined as
| Model | Epoch | Batch size | LR | Weight decay | # in-batch classes | Early stop | Training set | ||
| Train | Val | Train | Val | ||||||
| Pre-training stage 1 | |||||||||
| PEEB | 32 | 32 | 50 | 0.01 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 10 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 10 | Bird-11K | |
| Pre-training stage 2 | |||||||||
| PEEB | 32 | 32 | 50 | 0.01 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 5 | Bird-11K | |
| PEEB | 32 | 32 | 50 | 0.001 | 48 | 50 | 5 | Bird-11K | |
| Model | Fine-tune from | Epoch | Batch size | LR | Weight decay | Early stop | Training set |
| PEEB | PEEB | 30 | 32 | 0.001 | 5 | CUB | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB ZSL (2015) | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB-SCS | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | CUB-SCE | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | NABirds-SCS | |
| PEEB | PEEB | 5 | 32 | 0.001 | 5 | NABirds-SCE |
Appendix B Model and dataset notations
B.1 Dataset notations
Following the conventional setup of ZSL, we execute certain exclusions to make sure none of the test classes or descriptors are exposed during pre-training. That is, Bird-11K and Bird-11K exclude all CUB and NABirds classes, respectively. For GZSL, we exclude all test sets in CUB, NABirds, and iNaturalist, denoted as Bird-11K. We provide detailed statistics for the three pre-training sets in Table A3.
| Training set | Number of images | Number of classes | ||
| Train | Val | Train | Val | |
| Bird-11K | 234,693 | 29,234 | 10,740 | 9,746 |
| Bird-11K | 244,182 | 28,824 | 10,602 | 9,608 |
| Bird-11K | 216,588 | 27,996 | 10,326 | 9,332 |
B.2 Model notations
We adopt a strategy based on the datasets excluded during training to simplify our model naming convention. Specifically:
PEEB is pre-trained model using Bird-11K datset.
PEEB is pre-trained model using the Bird-11K dataset.
PEEB is pre-trained model using the Bird-11K dataset.
We named finetuned models after the pre-trained model and the finetuned training set. For example, PEEB is finetuned from PEEB, on CUB training set.
Appendix C Generating part-based descriptors
CUB annotations initially comprise 15 bird parts. However, distinctions between the left and right part are not essential to our method, we merge them into a single part (i.e., “left-wing” and “right-wing” are merged into “wings”) Hence, we distilled the original setup into 12 definitive parts: back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail, throat. To compile visual part-based descriptors for all bird species within Bird-11K, we prompted GPT-4 (OpenAI, 2023) with the following input template:
A bird has 12 parts: back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail and throat. Visually describe all parts of {class name} bird in a short phrase in bullet points using the format ‘part: short phrase’
Where {class name} is substituted for a given bird name (e.g., Painted Bunting).
The output is a set of twelve descriptors corresponding to twelve parts of the query species. e.g. The response for Cardinal is:
Appendix D Datasets
D.1
Bird-11K
We provide a brief statistic of Bird-11K in Table A4. Bird-11K is a diverse and long-tailed bird-image dataset. The descriptors generated by GPT-4 are in English and only describe the visual features of the corresponding class. We propose Bird-11K for academic research only.
| Dataset | # of Images | # of Species |
| CUB-200-2011 (Wah et al., 2011) | 12,000 | 200 |
| Indian Birds (Vaibhav Rokde, 2023) | 37,000 | 25 |
| NABirds v1 (Van Horn et al., 2015) | 48,000 | 400 |
| Birdsnap v7 (Berg et al., 2014) | 49,829 | 500 |
| iNaturalist 2021-birds (Van Horn et al., 2021) | 74,300 | 1,320 |
| ImageNet-birds (Deng et al., 2009) | 76,700 | 59 |
| BIRDS 525 (Piosenka, 2022) | 89,885 | 525 |
| Macaulay Library at the Cornell Lab of Ornithology* | 55,283 | 10,534 |
| Bird-11K (Raw Data) | 440,934 | 11,097 |
| Bird-11K (pre-training set) | 294,528 | 10,811 |
Data splits
We provide data splits and metadata, e.g., file names, image size, and bounding boxes, along with the instruction of Bird-11K construction in our repository. Note that the Bird-11K dataset is for pre-training purposes; it is important to execute exclusion based on the test set.
License and terms
CUB (Wah et al., 2011): The dataset can be freely used for academic and research purposes; commercial use is restricted.
Indian Birds (Vaibhav Rokde, 2023): CC0: Public Domain.
NABirds-v1 (Van Horn et al., 2015): For non-commercial research purposes, other use is restricted 333See Terms of Use here for detail: .
Birdsnap-v7 (Berg et al., 2014): The dataset creator provides no specific license or terms of use. We only use this dataset for academic research until more specific details can be obtained.
ImageNet-birds (Deng et al., 2009): BSD-3-Clause license.
BIRDS 525 (Piosenka, 2022): CC0: Public Domain
Cornell eBird: We used the following 55,384 recordings from the Macaulay Library at the Cornell Lab of Ornithology. The data is for academic and research purposes only, not publicly accessible unless requested. (Please refer to our Supplementary Material for the full list):
ML187387391, ML187387411, ML187387421, ML187387431, ML262407521, ML262407481, ML262407531, ML262407491, ML262407511, ML257194111 ML257194071, ML257194081, ML257194061, ML495670791, ML495670781, ML495670801, ML495670771, ML183436431, ML183436451, ML183436441 ML183436411, ML183436421, ML256545901, ML256545891, ML256545841, ML256545851, ML256545831, ML169637941, ML238083081, ML169637881 ML169637911, ML238083111, ML238083051, ML169637971, ML299670841, ML64989231, ML299670831, ML64989241, ML299670791, ML64989251 ML246866001, ML246865941, ML246866011, ML246865961, ML246865971, ML333411961, ML240835531, ML240835541, ML240835701, ML240835591 ML245260391, ML245260341, ML245260371, ML245260411, ML245260421, ML245260431, ML245260441, ML240866351, ML240866331, ML240866321 ML240866341, ML240866371, ML248318661, ML248318571, ML248318591, ML248318581, ML248318631, ML245204281, ML245204311, ML245204371 ML245204381, ML245204291, ML245603571, ML245603521, ML245603511, ML245603491, ML245603501, ML245603601, ML245257771, ML245257651 ML245257631, ML245257661, ML245257761, ML247221051, ML247221061, ML247221071, ML247221081, ML240365811, ML240365751, ML240365781 ML240365761, ML300579541, ML247298551, ML247298541, ML247298561, ML247298611, ML247298571, ML247298591, ML247298601, ML247298631…
D.2
Dog-140
To pre-train PEEB on dogs, we construct Dog-140 by combining dog images from ImageNet-21K and Stanford Dogs-120. Specifically, we selected 189 dog classes from ImageNet-21K, and based on Fédération Cynologique Internationale (FCI) (Fédération Cynologique Internationale (2023), FCI), we merged them with 120 classes from Stanford Dogs, ending up with 140 classes. After merging, Dog-140 has 206,076 images in total. We provide a class distribution analysis in Fig. A4, where we can find that Dog-140 is roughly class-balanced.
Data splits
Similar to Bird-11K, we provide data splits and metadata, e.g., file names, image size, and bounding boxes, along with the instruction of Dog-140 construction in our repository.
License and terms
Appendix E Additional results
E.1 PEEB outperforms M&V in CUB and NABirds in ZSL setting
To rigorously evaluate the ZSL capabilities of our pre-trained models, we introduce a stress test on the CUB and NABirds datasets. The crux of this test involves excluding all classes from the target dataset (CUB or NABirds) during the pre-training. The exclusion ensures that the model has no prior exposure to these classes. Subsequently, we measure the classification accuracy on the target dataset, comparing our results against benchmarks set by CLIP and M&V in the scientific name test. In this experiment, we consider the scientific name test a ZSL test for CLIP and use them as the baseline because the frequencies of scientific names are much lower than common ones.
Experiment
To conduct this test, we pre-train our model on Bird-11K and Bird-11K, which deliberately exclude images bearing the same class label as the target dataset. Specifically, we test on our pre-train model PEEB and PEEB (see Table A1 for details), respectively.
Results
The primary objective is to ascertain the superiority of our pre-trained model, PEEB, against benchmarks like CLIP and M&V. For CUB, our method reported a classification accuracy of 17.9%, contrasting the 5.95% and 7.66% achieved by CLIP and M&V, respectively, as shown in Table A5. The PEEB score, which is substantially higher (+10) than M&V, highlights the advantages of our part-based classification. On NABirds, our method surpasses CLIP and M&V by +1 point. The performance disparity between CUB and NABirds can be attributed to two factors: The elevated complexity of the task (555-way classification for NABirds versus 200-way for CUB) and the marked reduction in training data. An auxiliary observation, detailed in Sec. E.3, indicates that our pre-trained model necessitates at least 250k images to achieve admirable classification accuracy on CUB, but we only have 210k images training images in Bird-11K (the variants of Bird-11K with classes excluded for ZSL testing are described in Table A3).
| Method | CLIP | M&V | PEEB (ours) |
| CUB | 5.95 | 7.66 | 17.90 |
| NABirds | 4.73 | 6.27 | 7.47 |
E.2 Performance measurement on different noisy levels
In our evaluations, as indicated in Table 2, we discerned a marked performance disparity between the iNaturalist dataset and others. Probing this further, we identified image noise as a principal contributor to these discrepancies.
Experiment
A qualitative assessment of the iNaturalist test images revealed a significantly higher noise level than CUB or NABirds. To systematically study this, we utilize the object detector OWL-ViT to measure the size of the bird within the images. We formulated two filtered test sets based on the detector’s output, categorizing them by the bird’s size, specifically, the detected bounding box. Images were filtered out if the bird’s size did not exceed predetermined thresholds (areas of or pixels). Larger birds naturally reduced other content by occupying more image space, thus serving as a proxy for reduced noise. All three test sets, including the original, were evaluated using our pre-trained model PEEB.
Results
The results presented in Table A6 reveal a clear trend: as the image noise level decreases, the classification accuracy consistently improves, with gains ranging from (+6 to +17) points across the various methods. Notably, cleaner images consistently yield better results. At each noise level, our method outperforms the alternatives. While our method exhibits an impressive (+17 points) accuracy boost on the cleanest test set, this substantial gain also indicates that our model is sensitive to image noise.
| Splits | CLIP | M&V | PEEB (ours) |
| Original | 16.36 | 17.57 | 25.74 |
| pixels | 20.18 | 21.66 | 35.32 |
| pixels | 22.88 | 24.90 | 42.55 |
E.3 Number of training images is the most critical factor towards classification accuracy
Bird-11K, as shown in Fig. 5(a), is a highly imbalanced dataset characterized by a large amount of long-tailed classes. We conduct a comprehensive study to discern how variations in the number of classes and images affect the classification accuracy of our pre-trained models. Predictably, the volume of training images occurred as the most influential factor. However, a noteworthy observation was that the abundance of long-tailed data enhanced the model’s accuracy by approximately +1.5 points.
Experiment
We curated eight training sets based on varying class counts: 200, 500, 1,000, 2,000, 4,000, 6,000, 8,000, and 10,740. For each set, we maximized the number of training images. It is important to note that a set with a lesser class count is inherently a subset of one with a higher count. For instance, the 500-class set is a subset of the 2,000-class set. For each split, we apply the same training strategy as in Sec. 4.3.1, and choose the checkpoint with the best validation accuracy. We consider the CUB test set as a generic testing benchmark for all variants.
Results
As illustrated in Figure Fig. 5(b), there is a pronounced correlation between the increase in the number of images and the corresponding surge in accuracy. For instance, an increment from 106K to 164K images led to a rise in classification accuracy from 30.05% to 43.11%. The accuracy appears to stabilize around 60% when the image count approaches 250K. This trend strongly suggests that the volume of training images is the most critical factor for the pre-trained model. We believe that the accuracy of the pre-trained model could be further enhanced if enough data is provided. Interestingly, a substantial amount of long-tailed data bolsters the model’s performance, evident from +1.5 points accuracy improvement when comparing models trained on 2,000 classes to those on 10,740 classes. Note that the additional classes in the latter set averaged merely 2.2 images per class.
E.4 Ablation study on the influence of parts utilized
In this ablation study, we aimed to measure the impact of varying the number of distinct “parts” (back, beak, belly, breast, crown, forehead, eyes, legs, wings, nape, tail, and throat) used in our model. We experiment with a range from a single part to all 12 identifiable parts. Interestingly, even with a solitary part, the model could make correct predictions, though there was an evident decline in performance, approximately -20 points.
Experiment
Our testing ground is the pre-trained model PEEB, evaluated against the CUB test set. We assessed the model’s prowess utilizing various subsets of parts: 1, 3, 5, 8, and all 12. These subsets were derived based on the frequency of visibility of the parts within the CUB dataset, enabling us to compare the model’s performance when relying on the most frequently visible parts versus the least. For comparison, we also conduct a similar experiment on M&V, where we only use 1, 3, 5, 8, and 12 descriptors (if possible).
Results
Relying solely on the most frequent part led to a decline in classification accuracy by around -20 points, registering at 45.44% (Table A7). In contrast, utilizing the least frequent part resulted in a sharper drop of around -27, with an accuracy of 37.02%. As the model was furnished with increasing parts, its accuracy improved incrementally. The data underscores that optimal performance, an accuracy of 64.33%, is attained when all 12 parts are included. For M&V, the accuracy keeps increasing homogeneously from 5 to 12 descriptors, hinting that accuracy may increase further by increasing the number of descriptors.
| Number of Parts (descriptors) | 1 | 3 | 5 | 8 | 12 |
| Accuracy (most frequent parts) | 45.44 | 56.48 | 59.89 | 61.32 | 64.33 |
| Accuracy (least frequent parts) | 37.02 | 55.51 | 60.04 | 61.13 | 64.33 |
| Accuracy of Menon and Vondrick (2023) | 51.93 | 52.87 | 52.83 | 53.33 | 53.92 |
E.5 Training is essential for PEEB’s classification efficacy
In this ablation study, we highlight the pivotal role of training in the performance of PEEB on bird classification tasks. We demonstrate that without adequate tuning, the results are indistinguishable from random chance.
Experiment
We conduct the experiment based on OWL-ViT. We retain all components as illustrated in Fig. A1, with one exception: we substitute the Part MLP with the MLP layer present in the box prediction head of OWL-ViT because the proposed layers require training. The MLP layers in the box prediction head project the part embeddings to match the dimensionality of the text embeddings. Our focus is on assessing the classification accuracy of the untuned PEEB on two datasets: CUB and NABirds.
Results
Table A8 reveals the outcomes of our experiment. Without training, PEEB yields classification accuracies of 0.55% for CUB and 0.31% for NABirds, both of which are proximate to random chance (0.5% for CUB and 0.1% for NABirds). However, with training, the model’s performance dramatically transforms: 64.33% for CUB (an increase of +63.78 points) and 69.03% for NABirds (a leap of +68.72 points) for PEEB. These pronounced disparities underscore the vital role of training in PEEB.
| CUB | NABirds | |
| PEEB (no training) | 0.55 | 0.31 |
| PEEB pre-trained | 64.33 | 69.03 |
| PEEB finetuned | 86.73 | - |
E.6 Failure analysis
Since PEEB has two branches, box detection, and descriptor matching, we would like to find out, in the failure case, what is the main cause. i.e., is it because of the mismatch in the descriptor to the part embeddings? Or is it because the box detection is wrong? From our ablation study, it turns out that most errors come from the descriptor-part matching.
Experiment
We conduct the experiment with PEEB on CUB test set. Specifically, we measure the box detection accuracy based on the key point annotation in CUB dataset, i.e., We consider the box prediction as correct if the prediction includes the human-annotated key point. We report the box prediction error rate (in %) based on parts.
Results
As shown in Table A9, the average error rate difference between success and failure cases is merely 0.38. That is, in terms of box prediction, the accuracy is almost the same, disregarding the correctness of bird identification. It indicates that the prediction error is predominantly due to the mismatch between descriptors and part embeddings. We also noted that some parts, like Nape and Throat, have a very high average error rate, which may greatly increase the matching difficulties between descriptors and part embeddings.
| Body Part | Average | Back | Beak | Belly | Breast | Crown | Forehead | Eyes | Legs | Wings | Nape | Tail | Throat |
| Failure Cases | 16.52 | 23.38 | 3.28 | 8.06 | 15.96 | 7.41 | 24.72 | 7.29 | 5.63 | 3.36 | 64.79 | 7.25 | 27.07 |
| Success Cases | 16.14 | 23.03 | 2.96 | 7.44 | 18.64 | 7.13 | 21.53 | 3.93 | 6.85 | 2.68 | 68.66 | 6.40 | 24.38 |
| Difference | 0.38 | 0.35 | 0.33 | 0.62 | -2.68 | 0.28 | 3.19 | 3.36 | -1.22 | 0.68 | -3.87 | 0.85 | 2.68 |
E.7 Evaluation of predicted boxes from PEEB
Our proposed method primarily aims to facilitate part-based classification. While the core objective is not object detection, retaining the box prediction component is paramount for ensuring model explainability. This section delves into an evaluation of the box prediction performance of our method against the OWL-ViT model.
Experiment
Given our focus on part-based classification, we aimed to ascertain the quality of our model’s box predictions. To this end, we employed two metrics: mean Intersection over Union (IoU) and precision based on key points. We opted for mean IoU over the conventional mAP because: (1) Ground-truth boxes for bird parts are absent, and (2) our model is constrained to predict a single box per part, ensuring a recall of one. Thus, we treat OWL-ViT’s boxes as the ground truth and evaluate the box overlap through mean IoU. Furthermore, leveraging human-annotated key points for bird parts, we measure the precision of predicted boxes by determining if they contain the corresponding key points. We evaluate our finetuned models on their corresponding test sets. For instance, PEEB, finetuned based on the CUB split (Akata et al., 2015), is evaluated on the CUB test set.
Results
Our evaluation, as presented in Table A10, shows that PEEB’s box predictions do not match those of OWL-ViT. Specifically, on average, there is a -5 to -10 points reduction in mean IoU for CUB and NABirds datasets, respectively. The disparity is less distinct when examining precision based on human-annotated key points; our method records about -0.14 points lower precision for CUB and -3.17 points for NABirds compared to those for OWL-ViT. These observations reinforce that while PEEB’s box predictions might not rival these dedicated object detection models, they consistently highlight the same parts identified by such models as shown in Fig. A6. It is important to note that our approach utilized the same visual embeddings for both classification and box prediction tasks. This alignment emphasizes the part-based nature of our model’s predictions.
| Models | Mean IoU | |||
| (1) All | (2) w/ Keypoints | Precision | ||
| CUB | OWL-ViT | 100.00 | 100.00 | 83.83 |
| OWL-ViT | 44.41 | 49.65 | 83.53 | |
| PEEB (Average) | 35.98 | 40.14 | 83.39 | |
| PEEB | 37.45 | 41.79 | 81.55 | |
| PEEB | 35.11 | 39.14 | 82.72 | |
| PEEB | 35.77 | 39.96 | 84.89 | |
| PEEB | 35.58 | 39.67 | 84.38 | |
| NABirds | OWL-ViT | 100.00 | 100.00 | 85.01 |
| OWL-ViT | 40.14 | 47.63 | 83.89 | |
| PEEB (Average) | 36.47 | 42.01 | 80.72 | |
| PEEB | 36.45 | 42.03 | 80.09 | |
| PEEB | 36.49 | 41.99 | 81.34 | |
Appendix F Study on GPT-4 generated descriptors
F.1 Assessment of the generated part-based descriptors
We test GPT-4V on the CUB test set using the generated descriptors of 200 classes to assess their usability. Specifically, we feed GPT-4V with each test image encoded in the payload and 200 sets of part-based descriptors through a carefully designed prompt (Table A11). GPT-4V is asked to output one of 200 provided class names to compute the classification accuracy. As a result, GPT-4V achieves 69.4% accuracy which is slightly higher than PEEB’s generalized zero-shot accuracy (64.33%) and significantly lower than PEEB results after finetuning (86-88%).
F.2 Noise measurement in GPT-4 generated descriptors
In this section, we conduct an empirical analysis to quantify the noise in descriptors generated by GPT-4 for 20 different classes within the CUB dataset. To achieve this, we manually inspect each descriptor and tally the instances where at least one factual error is present. Our findings reveal that every one of the 20 classes contains descriptors with errors, and on average, 45% of the descriptors necessitate corrections. This substantial noise level underscores the need for further refinement in our work, particularly in text descriptors.
We observe a notably high error rate in descriptors on the back and wings, with approximately 60% of these containing inaccurate information (refer to Table A12). This could be attributed to the challenges in distinguishing between the back and wings, given that the back is typically positioned behind the wings, yet exhibits considerable variability in size and shape. Addressing all descriptor issues by revising all 11,000 fine-grained descriptors would demand a significant investment of time and resources, which is beyond the scope of the current work. As such, we identify this as an area for future research and development, aiming to enhance the quality of the Bird-11K dataset.
| Back | Beak | Belly | Breast | Crown | Forehead | Eyes | Legs | Wings | Nape | Tail | Throat | Average | |
| Error Rate | 60 | 30 | 50 | 40 | 50 | 55 | 50 | 20 | 60 | 50 | 35 | 40 | 45 |
F.3 Revising descriptors improves classification accuracy
As mentioned in the limitation section, the descriptors are generated from GPT-4 and therefore noisy and incorrect. Given that PEEB accepts open vocabulary inputs for classification, a natural way to improve classification accuracy is to improve the correctness of the descriptors.
Experiment
We first collect descriptors of 183 CUB classes from AllAboutBirds. We then prompt GPT-4 to revise our original descriptors by providing the collected descriptor. We revise the descriptors with the following prompt:
Given the following descriptors of {class name}: {AllAboutBirds descriptors}. Can you revise the incorrect items below (if any) of this bird, return them as a Python dictionary, and use the key as the part name for each item? If a partś descriptor is not specifically described or cannot be inferred from the definition, use your own knowledge. Otherwise, leave as is. Note: please use a double quotation mark for each item such that it works with JSON format.
{Original descriptors}
Where {class name} the placeholder for the class name, {AllAboutBirds descriptors} is the description collected from AllAboutBirds, {Original descriptors} is the descriptors we used for training.
Due to the errors in the descriptors we used to train PEEB, simply replacing the descriptors with their revised version does not lead to better performance. Because the incorrect descriptors in training change the meaning of some of the phrases. For example, the belly of Blue bunting is pure blue, but the descriptors from GPT-4 is soft, creamy white. In addition, the GPT-4 uses the exact same descriptor in the belly for other classes, e.g., Blue breasted quail, which should be cinnamon. Blue Fronted Flycatcher, which should be yellow. Training the same descriptors with different colors confuses the model, and the model will convey the phrase “creamy white” with a different meaning to humans. Therefore, simply changing the descriptors to their’ revised version will not work. We empirically inspect the descriptors that PEEB can correctly respond to and replace the class descriptors with the revised version. Specifically, we replace the descriptors of 17 classes in CUB and test the classification accuracy on PEEB.
Results
As shown in Table A13, the overall accuracy increases by +0.8 points.
The average improvement of the revised class is around +10.8, hitting that if we have correct descriptors of all classes, we may significantly improve the classification accuracy of the pre-trained model. However, correcting all 11k class descriptors is too expensive and out of the scope of this work. We leave it as a further direction of improving the part-based bird classification.
| Descriptors | Original | Partially Revised | Avg. Improvement |
| PEEB | 64.33 | 65.14 | 10.80 |
Appendix G Qualitative Inspections
G.1 Visual comparison of predicted boxes
We provide a visual comparison of the box prediction from OWL-ViT, OWL-ViT, and PEEB in Fig. A6. We find that despite the fact that our predicted boxes have lower mean IoU compared to OWL-ViT, they are visually similar to the boxes as OWL-ViT.
OriginalPEEBOWL-ViTOWL-ViT
G.2 Qualitative examples of using randomized descriptors
We visually compare M&V and PEEB based on their utilization of descriptors. (Figs. A7, A8 and A9). Specifically, we randomly swap the descriptors of the classes and then use these randomized descriptors as textual inputs to the tested models to see how they perform. We observe that the scores from M&V tend to cluster closely together. Surprisingly, M&V’s prediction remains unchanged despite the inaccurate descriptors. In contrast, PEEB, when presented with randomized descriptors, attempts to identify the best match grounded on the given descriptors.
Original DescriptorRandom nonsense Descriptor
M&V
PEEB
Original descriptorRandom nonsense descriptor
M&V
PEEB
Original descriptorRandom nonsense descriptor
M&V
PEEB
G.3 Examples of PEEB explanations for birds
Figs. A10, A11 and A12 are examples of how PEEB makes classification based on the descriptors and how it can reject the predictions made by M&V. Since we aggregate all descriptors for the final decision, even if some of them are similar in two classes, our method can still differentiate them from other descriptors. For instance, in Fig. A10, while other descriptors are similar, PEEB can still reject chesnut-sided warbler thanks to the distinct features of forehead, throat and belly.
G.4 Examples of PEEB explanations for dogs
Figs. A13, A14 and A15 are examples of how PEEB makes classification based on the descriptors in Stanford Dogs dataset. We demonstrate that our model works well on dogs, which indicates that our proposed method is transferable to other domains while maintaining high-quality explainability as in birds.